### v1

# k* Distribution: Evaluating the Latent Space of Deep Neural Networks using Local Neighborhood Analysis

Abstract
Most examinations of neural networks’ learned latent spaces typically employ dimensionality reduction techniques such as t-SNE or UMAP. While these methods effectively capture the overall sample distribution in the entire learned latent space, they tend to distort the structure of sample distributions within specific classes in the subset of the latent space. This distortion complicates the task of easily distinguishing classes identifiable by neural networks. In response to this challenge, we introduce the k* Distribution methodology. This approach focuses on capturing the characteristics and structure of sample distributions for individual classes within the subset of the learned latent space using local neighborhood analysis. The key concept is to facilitate easy comparison of different k* distributions, enabling analysis of how various classes are processed by the same neural network. This provides a more profound understanding of existing contemporary visualizations. Our study reveals three distinct distributions of samples within the learned latent space subset: a) Fractured, b) Overlapped, and c) Clustered. We note and demonstrate that the distribution of samples within the network’s learned latent space significantly varies depending on the class. Furthermore, we illustrate that our analysis can be applied to explore the latent space of diverse neural network architectures, various layers within neural networks, transformations applied to input samples, and the distribution of training and testing data for neural networks. We anticipate that our approach will facilitate more targeted investigations into neural networks by collectively examining the distribution of different samples within the learned latent space.
Index Terms:
Deep Learning, Convolution Neural Networks, Image Classification, Latent Space Visualization, Local Neighborhood Analysis
**footnotetext: These authors contributed equally to this work
I Introduction
Figure 1: Overview of three distinct basic patterns of k* Distribution. Here we define the k* value of a sample point as ${}^{\text{th}}$-closest neighbor, which differs in class compared to the test point, i.e., the neighbor (sample) which breaks homogeneity in the local neighborhood of the test point. Pattern A (🟊normal-$\bigstar$) which has negatively skewed k* distribution (skewed towards high k* value) representing an ‘Fractured’ distribution of samples in latent space; Pattern B (normal-$\clubsuit$) which has almost uniform k* distribution representing a ‘Overlapped’ distribution of samples in latent space; Pattern C (normal-$\spadesuit$) which has positively skewed k* distribution (skewed towards low k* value) representing a ‘Clustered’ distribution of samples in latent space.
Figure 2: Overview of the framework to create k* Distribution. We use the learned features of a neural network to compute k* values of individual evaluated sample and then compute the k* distribution for a particular class.
A significant portion of neural network research relies on creating tools to comprehend the acquired latent space and unveil the inner workings of neural networks, which are often viewed as black boxes. Nevertheless, if researchers are equipped with the essential tools to grasp the learned latent space, it becomes feasible to delve deeper, uncovering insights and reasons that can guide further research in the field of neural networks. Analyzing the neural network’s latent space poses challenges due to its intricate non-convex characteristics. However, directly evaluating and comparing the configuration and distribution of data within high-dimensional latent spaces persists despite the development of various visualization tools.
Existing tools often rely on dimensionality reduction techniques like tSNE [1] and UMAP [2] to create $2$ or $3$-dimensional scatter plots for individual latent spaces, offering a broad overview. They rely on dimensionality reduction methods rooted in manifold learning which take into account specific manifold characteristics, preserving them while modifying other factors and attributes.
Analyzing the configurations and structures of various sample distributions associated with a specific class within the latent space presents a challenge using existing visualization methods. This challenge extends to the comparison of multiple sample distributions which limits the analysis of the latent spaces. Additionally, visually comparing multiple latent spaces side by side can be overwhelming and confusing, especially when dealing with numerous points and varying degrees of distortion [3, 4, 5, 6, 7].
If there is too much information kept, it is hard to understand. At the same time, if the information is filtered enough, there is always some perspective lacking. Therefore, some complementary tools might be crucial in interpreting these complex latent spaces.
This article introduces such a tool designed to enhance the understanding of the distribution of samples associated with ‘classes/features’ in the learned latent space generated by neural networks. The proposed approach involves leveraging the local neighborhood relationships among these samples in the latent space. More precisely, the nearest-neighbor method is applied to the learned latent space to identify a k*-nearest neighbor (sample) within the local neighborhood, featuring a different class from the test sample. This process evaluates the disruption of the homogeneity within the local neighborhood of the tested sample. Subsequently, a distribution of k*—referred to as the k* distribution—is generated, incorporating all the samples belonging to a given class. Through an analysis of the k* distribution, three distinct patterns of distribution of samples in the learned latent space are identified as illustrated in Figure 1:
Pattern A (🟊normal-$\bigstar$) representing Fractured distribution of samples in latent space,
Pattern B (normal-$\clubsuit$) representing Overlapped distribution of samples in latent space, or
Pattern C (normal-$\spadesuit$) representing Clustered distribution of samples in latent space.
Contributions: This article provides,
[…]
A new interpretation of latent space learned by the neural network, relying on local neighborhood relationships and homogeneity.
[…]
Identification of various distribution patterns of samples in the latent space based on neighborhood characteristics (see Figure 1).
[…]
A model-agnostic latent space analysis of neural networks, focusing on samples from a single class (see Figure 2).
[…]
A method for straightforwardly comparing different classes and understanding how samples from various classes are distributed in the learned latent space (see Figure 3).
II Related Works
[…]
II-A Visualising the Latent Space of Neural Networks using Dimensionality Reduction Techniques:
Given that neural networks operate in a high-dimensional space, visualizing their latent space directly is challenging. To overcome this limitation, various researchers employ dimensionality reduction techniques to represent the latent space in 2 or 3 dimensions. There are a lot of techniques for visualizing the latent space like, [1, 2, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25].
Many of these algorithms require hyperparameter tuning for optimal visualization, making it challenging to ensure a fair comparison between latent spaces of varying dimensions. Additionally, while these visualizations effectively capture some perspective on the structure of the learned latent space, using them to draw a clear comparison between different local structures is challenging [3, 4, 5, 6, 7].
The effectiveness of dimensionality reduction techniques becomes apparent in instances where the latent space is well-organized and has successfully assimilated the intended information. In such cases, these visualizations demonstrate utility by aligning with pre-established interpretations. Conversely, when applied to latent space lacking clearly defined pre-established knowledge, the efficacy diminishes for such visualizations where the lack of known structure leads to a ‘blob of points’ in the visualization [7].
II-B Visualising Association of Features in the Latent Space of Neural Network:
Various approaches employ visualizations to understand how neural networks interact with features. Analyzing the responses of units in hidden layers to features provides insights into the learned latent space, elucidating which features the network prioritizes [26, 27]. This involves evaluating individual unit responses or combinations of specific inputs [28].
Another visualization method focuses on understanding neural network attention, revealing which parts of an image are emphasized. Salient regions in the input that influence the network’s output are identified using saliency maps [29] and gradient-weighted class-activation maps [30]. These maps quantify conspicuity at each visual field location, guiding the selection of attended locations based on saliency distribution.
Additionally, interpretable concepts, defined as groups of latent variables in the space that are meaningful, may manifest in the latent space, further contributing to our understanding [26, 27, 31, 32, 33]. In the domain of natural language processing, these interpretations often provide insight into learning semantic and linguistic concepts provided by the lexical object [34, 35, 36]. These interpretations are often effectively visualized using analytic systems like [37, 38, 39, 40, 41, 42, 43], however, the challenge to draw a clear comparison between different local structures and multiple latent spaces persists.
To examine the distribution of samples in the neighborhood of a specific sample and analyze the distribution of samples sharing a predefined label, we suggest employing the k* distribution. With our proposed approach, we present a framework (see Figure 2) for effectively comparing various latent spaces and sub-spaces with diverse sample distributions. This framework enables meaningful comparisons between the distribution of samples belonging to different classes and multiple latent spaces, offering additional insights into latent spaces beyond what existing visualizations provide.
III k* Distribution: Analysing Homogeneity in the Local Neighborhood of Samples
We propose a methodology based on the local neighborhood to analyze the hyper-dimensional latent space learned by neural networks (see Figure 2). The distribution of samples and clusters in the learned latent space is analyzed by associating them with classes. Through this approach, we gain insights into the distribution patterns (see Figure 1) within the learned latent space and identify the clusters formed, thereby enhancing our comprehension of the latent space.
Visualizations inevitably involve trade-offs as they strive to make information comprehensible from specific perspectives. The concept of nearest neighbors is arguably the most relevant when portraying the structure of the distribution of samples, like clustering or fragmentation in a latent space. Alternative measures like the absolute distance can offer insight into other facets of the latent space, but they may not effectively reveal the intricate high-dimensional fracturing. Further projections of high dimensional latent space using these distances can obscure the crucial information about neighboring samples.
The Nearest Neighbours method, a widely recognized non-parametric technique, is employed enabling us to understand the positioning of latent variables in the vicinity of a latent variable in space and their corresponding class distribution. This method aids in gauging the relative distances between latent variables and grouping them into clusters. The underlying principle is grounded in the notion that local neighborhoods offer a reasonable estimate of the sample distribution within the latent space.
By exploring the concept of neighborhoods and their utility in validating the local relationship among features in latent space, we introduce the k* distribution in this context. The k* distribution is constructed by taking a latent space as input and analyzing the neighborhood of a sample as shown in Figure 3.
The measurement of the nearest neighbor index allows us to address the sparsity inherent in high-dimensional space. Rather than relying on absolute distance values between points, a relative measure through the concept of neighborhood is utilized. This approach facilitates the comparison of two distinct latent spaces with varying dimensionalities. Importantly, the neighborhood concept is dimensionality-independent, making this technique applicable and effective across both low and high dimensions.
The index of ${}^{\text{th}}$ nearest neighbor (sample) that belongs to a different class than the test sample disrupting the uniformity of the test sample’s neighborhood is quantified. This index is termed the k* value. Essentially, a high k* value indicates that the sample is surrounded by similar points that share the same class. In other words, a high k* value represents that a neighbor with a different class will be situated at a considerable distance from the neighborhood of the specified sample. In particular, the homogeneity of each class cluster within the learned latent space can be assessed using k* (see Figure 3). Additionally, one can ascertain whether a cluster remains cohesive for a given class or if it is fragmented across various spatial locations.
Let us consider a collection of sample-label pairs $X$: $(x_{1},Y_{1}),(x_{2},Y_{2}),...,(x_{n},Y_{n})$ where $x$ is the input sample and $Y$ is the label for the input sample. Here, the latent space is embedded with such $x$ points. Let $S$ be the set of all samples $x_{p}\in X$ such that they have the same label $c$, i.e.,
$S_{c}=\{x_{i}\mid\forall x_{i}\in X~{}\text{such that}~{}Y_{i}=c\}$ (1)
The distance $D$ of sample $x_{p}$ from all other samples can be formulated as follows:
$D(x_{p})=\{\text{distance}(x_{p},x_{i})\mid\forall(x_{i},y_{i})\in X\}$ (2)
where $\text{distance}(a,b)$ is the distance between two samples $a$ and $b$. The commonly employed distance function is the Minkowski distance, which is a generalization of various distance metrics, including Euclidean and Cityblock distances. The Minkowski distance of order $p$ between two points $a=(a_{1},a_{2}\ldots a_{r})$ and $b=(b_{1},b_{2}\ldots b_{r})$ in $r$ dimensional space is given by,
$\text{distance}(a,b)_{\text{Minkowski},p}=\left(\sum_{i=1}^{n}|a_{i}-b_{i}|^{r}\right)^{1/r}$ (3)
Here, when $p=2$, it represents Euclidean distance ($l_{2}$ norm); when $p=1$, it represents Cityblock distance ($l_{1}$ norm); and when $p=\infty$, it represents Maximum Norm distance ($l_{\infty}$ norm).
$i^{th}$ neighbour sample $x_{i}^{p}$ to sample $x_{p}$ is defined as,
$x_{i}^{p}=x_{k}\quad\text{where}\quad k\in\operatorname*{arg\,min}_{x_{k}\in P_{i}}~{}\text{distance}(x_{p},x_{k})$ (4)
such $\text{such that}\quad P_{i}=X-\{x_{j}^{p}\mid\forall j<i\}$
Similarly, local neighbourhood space $N_{p}$ of sample $(x_{p})$ can be defined as,
$N_{p}=(x_{0}^{p},x_{1}^{p}\ldots x_{n}^{p})$ (5)
[…]
$\text{k}^{*}_{p}=\operatorname*{arg\,min}_{(x_{p},Y_{p})}\{x_{i}^{p}\mid x_{i}^{p}\in N_{p},~{}Y_{i}^{p}\neq Y_{p}\}$ (6)
where $i$ is the index of the nearest neighbor, $Y_{p}$ is the label of test sample $x_{p}$ and $Y_{i}^{p}$ is the label of the nearest neighbor (sample) $x_{i}^{p}$ that differs compared to label $Y_{p}$. Then, the k* distribution $\text{k}^{*}(\cdot)$ of class $c$ can be defined as,
$\text{k}^{*}(S_{c})=\left\{\frac{\text{k}^{*}_{p}}{|S_{c}|}\mid\forall x_{p}\in S_{c}\right\}$ (7)
Here, $|S_{c}|$ is the cardinality of set $S_{c}$ representing the number of samples belonging to class $c$.
Based on the defined k* distribution we define certain metrics over it described below,
Mean of k* distribution
$\mu_{k^{*}}$ can be defined as,
$\mu_{k^{*}}=\frac{1}{|S|}\sum\text{k}^{*}(S)$ (8)
Standard Deviation of k* distribution
$\sigma_{k^{*}}$ can be defined as,
$\sigma_{k^{*}}=\left(\frac{1}{|S|}\sum(\text{k}^{*}(S)-\mu_{k^{*}})^{2}\right)^{\frac{1}{2}}$ (9)
measuring the amount of variation or dispersion in k* values.
Fischer-Pearson Skewness Coefficient of k* distribution
$\gamma_{k^{*}}$ can be defined as,
$\gamma_{k^{*}}=\frac{\frac{1}{|S|}\sum(\text{k}^{*}(S)-\mu_{k})^{3}}{\left(\frac{1}{|S|}\sum(\text{k}^{*}(S)-\mu_{k})^{2}\right)^{3/2}}$ (10)
It measures the asymmetry of the k* distribution about its mean $\mu_{k^{*}}$. The skewness coefficient can be positive, negative, or zero. A positive skewness indicates a distribution that is skewed to the left, i.e., towards lower k* metric values, while a negative skewness indicates a distribution that is skewed to the right, i.e., towards higher k* metric values.
Based on the k* distributions, we observe three distinct patterns (Figure 1) of sample distribution in the latent space:
Figure 4: Visualization of the distribution of samples in latent space using, (Left) k* distribution, and (Right) Dimentionality Reduction techniques like t-SNE (Top Left), Isomap (Top Right), PCA (Bottom Left), and UMAP (Bottom Right)  of all classes of 16-class-ImageNet for the Logit Layer of ResNet-50 Architecture. Note, how the distribution of samples for a particular class is easier to compare using k* distribution compared to using dimensionality reduction techniques.
Pattern A (🟊normal-$\bigstar$) Fractured distribution of samples:
[…]
Pattern B (normal-$\clubsuit$) Overlapped distribution of samples:
[…]
Pattern C (normal-$\spadesuit$) Clustered distribution of samples:
[…]
IV Experimental Results and Visualization
IV-A Common Experimental Setup
Datasets: Initially, we assess the k* distribution using the entire set of $1280$ images from the $16$-class-ImageNet dataset curated by Geirhos et al. [44]. This dataset contains $80$ images per each of the $16$ classes of the dataset. These are the $16$ entry-level categories from MS-COCO that have the highest number of ImageNet classes mapped via the WordNet hierarchy making them compatible with the $1,000$ classes of ImageNet-1k dataset [45]. The purpose of employing this dataset is to evaluate individual classes organized naturally.
In addition to the $16$-class-ImageNet dataset, we evaluate k* distribution comprehensively using $50,000$ validation images of the original ImageNet-1k dataset [45]. This dataset is divided into $50$ images for each of its $1,000$ classes of the dataset.
To further scrutinize the models’ latent spaces, we apply various transformations to the samples from both the $16$-class-ImageNet dataset and ImageNet-1k dataset. These transformations include Image Crop, the addition of Gaussian Noise, Rotation of the images, and the generation of Adversarial and Stylized Versions of the samples.
Moreover, to assess the k* distribution across various other tasks, we also use various other datasets for different tasks. For intent classification, we use the English subset of the MASSIVE [46] dataset. The testing samples contained $2,970$ samples split across $60$ distinct intent classes. For the keyword spotting task, we used the Speech Commands v0.02 dataset [47], which contains $4,890$ samples split across $36$ command categories. We utilize these datasets to analyze as described in Figure 2 and comprehend the characteristics associated with each pattern outlined in Figure 1.
Deep Neural Network Architectures: We analyze the latent space of pre-trained weights from various open-source neural networks referenced as we have used them in the experiments. Adversarial Attacks: We subject the trained ResNet-50 model to a Projected Gradient Descent (PGD) attack, as outlined in [48], employing a perturbation magnitude of ${\epsilon}=4/255$. This attack is aimed at gauging the extent of the effective alteration in representation induced by the adversarial samples.
Furthermore, we assess the model that has undergone adversarial training using the PGD attack. This evaluation aims to compare and contrast the latent space of the robust model with that of its non-robust counterpart.
TABLE I: Multi-category evaluation of various neural architectures by using statistical metrics across object categories. Metrics include Mean ($\mu_{k^{*}}$), Standard deviation ($\sigma_{k^{*}}$), Skewness Coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and prevailing Pattern (Pat) in the context of diverse visual patterns.
Airplane Bicycle Car
Architectures $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat
ResNet-50 [49] 0.57 0.25 -0.84 100.00 $\spadesuit$ 0.15 0.11 2.60 97.50 🟊$\bigstar$ 0.45 0.26 -0.17 93.75 $\clubsuit$
ResNeXt-101 [50] 0.54 0.23 -0.71 100.00 $\spadesuit$ 0.22 0.17 1.23 100.00 🟊$\bigstar$ 0.50 0.28 -0.32 97.50 $\clubsuit$
EfficientNet-B0 [51] 0.80 0.19 -3.03 100.00 $\spadesuit$ 0.44 0.20 -0.53 96.25 $\spadesuit$ 0.54 0.20 -0.93 96.25 $\spadesuit$
ViT [52] 0.89 0.12 -5.08 100.00 $\spadesuit$ 0.41 0.15 -0.20 98.75 $\clubsuit$ 0.52 0.20 -0.44 93.75 $\clubsuit$
Dog Elephant Truck
Architecture $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat
ResNet-50 [49] 0.14 0.11 2.88 95.00 🟊$\bigstar$ 0.62 0.29 -0.99 98.75 $\spadesuit$ 0.44 0.25 -0.20 98.75 $\clubsuit$
ResNeXt-101 [50] 0.14 0.12 2.52 97.50 🟊$\bigstar$ 0.71 0.27 -1.38 100.00 $\spadesuit$ 0.49 0.24 -0.45 98.75 $\clubsuit$
EfficientNet-B0 [51] 0.05 0.03 7.89 100.00 🟊$\bigstar$ 0.77 0.18 -2.83 100.00 $\spadesuit$ 0.48 0.20 -0.86 98.75 $\spadesuit$
ViT-B [52] 0.05 0.03 8.06 98.75 🟊$\bigstar$ 0.92 0.09 -5.24 100.00 $\spadesuit$ 0.53 0.19 -0.67 100.00 $\spadesuit$
TABLE II: Performance by various neural architectures across varied visual patterns - Pattern A (🟊normal-$\bigstar$), Pattern B (normal-$\clubsuit$), Pattern C (normal-$\spadesuit$), and $1,000$ ImageNet1k Classes. The table presents key statistics, including an average of Mean ($\mu_{k^{*}}$), Skewness coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and Number of Classes (N).
Average of Classes with Pattern A (🟊normal-$\bigstar$) (Fractured) Average of Classes with Pattern B (normal-$\clubsuit$) (Overlapped) Average of Classes with Pattern C (normal-$\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
Architecture $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc
ResNet-50 [49] 0.10 4.51 74.88 935 0.37 0.14 93.96 56 0.61 -1.04 97.11 9 0.12 4.21 76.15
ResNeXt-101 [50] 0.11 4.23 75.05 865 0.38 0.08 93.44 104 0.60 -1.12 96.06 31 0.15 3.63 77.62
EfficientNet-B0 [51] 0.10 4.56 74.66 840 0.37 0.01 92.50 105 0.59 -1.40 95.45 55 0.15 3.76 77.68
ViT [52] 0.12 3.83 74.33 604 0.35 0.01 87.78 153 0.64 -1.85 93.59 243 0.28 1.86 81.07
TABLE III: Multi-category evaluation of various layers of ResNet-50 by using statistical metrics across object categories. Metrics include Mean ($\mu_{k^{*}}$), Standard deviation ($\sigma_{k^{*}}$), Skewness Coefficient ($\gamma_{k^{*}}$), and prevailing Pattern (Pat) in the context of diverse visual patterns.
Airplane Bicycle Car
Layer Name $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat
Logit Layer 0.57 0.25 -0.84 $\spadesuit$ 0.15 0.11 2.60 🟊$\bigstar$ 0.45 0.26 -0.17 $\clubsuit$
Average Pooling 0.59 0.25 -0.93 $\clubsuit$ 0.15 0.12 2.59 🟊$\bigstar$ 0.42 0.26 -0.07 $\clubsuit$
Stage4 Block3 Conv3 0.27 0.16 0.76 $\clubsuit$ 0.08 0.06 5.50 🟊$\bigstar$ 0.06 0.06 6.12 🟊$\bigstar$
Stage3 Block6 Conv3 0.07 0.08 4.78 🟊$\bigstar$ 0.03 0.02 8.37 🟊$\bigstar$ 0.03 0.03 8.14 🟊$\bigstar$
Stage2 Block4 Conv3 0.01 0.00 8.88 🟊$\bigstar$ 0.01 0.00 8.89 🟊$\bigstar$ 0.01 0.00 8.89 🟊$\bigstar$
Dog Elephant Truck
Layer Name $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat
Logit Layer 0.14 0.11 2.88 🟊$\bigstar$ 0.62 0.29 -0.99 $\spadesuit$ 0.44 0.25 -0.20 $\clubsuit$
Average Pooling 0.09 0.07 4.98 🟊$\bigstar$ 0.62 0.28 -1.06 $\spadesuit$ 0.43 0.25 -0.19 $\clubsuit$
Stage4 Block3 Conv3 0.01 0.00 8.87 🟊$\bigstar$ 0.10 0.08 4.41 🟊$\bigstar$ 0.06 0.05 6.55 🟊$\bigstar$
Stage3 Block6 Conv3 0.02 0.01 8.82 🟊$\bigstar$ 0.03 0.02 8.24 🟊$\bigstar$ 0.02 0.01 8.71 🟊$\bigstar$
Stage2 Block4 Conv3 0.01 0.00 8.89 🟊$\bigstar$ 0.01 0.00 8.89 🟊$\bigstar$ 0.01 0.00 8.89 🟊$\bigstar$
TABLE IV: Performance by different layers of ResNet-50 across varied visual patterns - Pattern A (🟊normal-$\bigstar$), Pattern B (normal-$\clubsuit$), Pattern C (normal-$\spadesuit$), and $1,000$ ImageNet1k Classes. The table presents key statistics, including an average of Mean ($\mu_{k^{*}}$), Skewness coefficient ($\gamma_{k^{*}}$), and Number of Classes (N).
Average of Classes with Pattern A (🟊normal-$\bigstar$)  (Fractured) Average of Classes with Pattern B (normal-$\clubsuit$)  (Overlapped) Average of Classes with Pattern C (normal-$\spadesuit$)  (Clustered) Average of 1,000 ImageNet1k Classes
Layer Name $\mu_{k^{*}}$ $\gamma_{k^{*}}$ N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ N $\mu_{k^{*}}$ $\gamma_{k^{*}}$
Logit Layer 0.10 4.51 935 0.37 0.14 56 0.61 -1.04 9 0.12 4.21
Average Pooling 0.09 4.59 837 0.34 0.35 157 0.72 -1.73 6 0.13 3.89
Stage4 Block3 Conv3 0.05 5.90 998 0.33 0.48 2 0 0.05 5.89
Stage3 Block6 Conv3 0.02 6.96 1,000 0 0 0.02 6.96
Stage2 Block4 Conv3 0.02 7.00 1,000 0 0 0.02 7.00
TABLE V: Multi-category evaluation of models trained on different training distributions by using statistical metrics across object categories. Metrics include Mean ($\mu_{k^{*}}$), Standard deviation ($\sigma_{k^{*}}$), Skewness Coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and prevailing Pattern (Pat) in the context of diverse visual patterns.
Airplane Bicycle Car
Training Distribution $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat
ImageNet1k [45] 0.57 0.25 -0.84 100.00 $\spadesuit$ 0.15 0.11 2.60 97.50 🟊$\bigstar$ 0.45 0.26 -0.17 93.75 $\clubsuit$
Stylised ImageNet [53] 0.41 0.24 -0.29 95.00 $\clubsuit$ 0.13 0.09 3.72 93.75 🟊$\bigstar$ 0.23 0.18 1.12 87.50 🟊$\bigstar$
ImageNet1k + Stylised [53] 0.58 0.24 -0.97 100.00 $\spadesuit$ 0.17 0.13 1.97 98.75 🟊$\bigstar$ 0.43 0.27 -0.02 92.50 $\clubsuit$
Dog Elephant Truck
Training Distribution $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat
ImageNet1k [45] 0.14 0.11 2.88 95.00 🟊$\bigstar$ 0.62 0.29 -0.99 98.75 $\spadesuit$ 0.44 0.25 -0.20 98.75 $\clubsuit$
Stylised ImageNet [53] 0.09 0.07 5.20 95.00 🟊$\bigstar$ 0.41 0.28 0.01 98.75 $\clubsuit$ 0.26 0.21 0.75 97.50 🟊$\bigstar$
ImageNet1k + Stylised [53] 0.16 0.13 2.22 96.25 🟊$\bigstar$ 0.64 0.29 -0.85 100.00 $\spadesuit$ 0.40 0.22 -0.07 100.00 $\clubsuit$
TABLE VI: Performance by models trained on different training datasets across varied visual patterns - Pattern A (🟊normal-$\bigstar$), Pattern B (normal-$\clubsuit$), Pattern C (normal-$\spadesuit$), and $1,000$ ImageNet Classes. The table presents key statistics, including an average of Mean ($\mu_{k^{*}}$), Skewness coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and Number of Classes (N).
Average of Classes with Pattern A (🟊normal-$\bigstar$) (Fractured) Average of Classes with Pattern B (normal-$\clubsuit$) (Overlapped) Average of Classes with Pattern C (normal-$\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
Training Distribution $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc
ImageNet1k [45] 0.10 4.51 74.88 935 0.37 0.14 93.96 56 0.61 -1.04 97.11 9 0.12 4.21 76.15
Stylised ImageNet [53] 0.05 5.84 59.99 994 0.35 0.19 91.33 6 0 0.06 5.80 60.18
ImageNet1k + Stylised [53] 0.09 4.67 73.36 939 0.37 0.12 92.90 51 0.57 -0.98 96.20 10 0.11 4.38 74.59
TABLE VII: Multi-category evaluation of Robust (Adversarially Trained: AT) and Non-Robust (Standard Trained: ST) architectures by using statistical metrics across object categories. Metrics include Mean ($\mu_{k^{*}}$), Standard deviation ($\sigma_{k^{*}}$), Skewness Coefficient ($\gamma_{k^{*}}$), and prevailing Pattern (Pat) in the context of diverse visual patterns.
Airplane Bicycle Car
Architecture Type $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat
ResNet-50 [49] AT [54] 0.40 0.22 -0.40 $\clubsuit$ 0.13 0.09 3.56 🟊$\bigstar$ 0.35 0.23 0.18 $\clubsuit$
ST [49] 0.57 0.25 -0.84 $\spadesuit$ 0.15 0.11 2.60 🟊$\bigstar$ 0.45 0.26 -0.17 $\clubsuit$
WideResNet-50 [55] AT [54] 0.64 0.26 -1.51 $\spadesuit$ 0.16 0.10 2.76 🟊$\bigstar$ 0.57 0.27 -0.81 $\spadesuit$
ST [55] 0.87 0.17 -3.75 $\spadesuit$ 0.21 0.14 1.30 🟊$\bigstar$ 0.57 0.20 -0.57 $\spadesuit$
ViT-B [52] AT [54] 0.76 0.17 -3.26 $\spadesuit$ 0.25 0.16 0.84 🟊$\bigstar$ 0.62 0.23 -1.26 $\spadesuit$
ST [52] 0.91 0.19 -3.27 $\spadesuit$ 0.30 0.17 0.25 $\clubsuit$ 0.56 0.26 -0.75 $\spadesuit$
Dog Elephant Truck
Architecture Type $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat
ResNet-50 [49] AT [54] 0.06 0.05 6.34 🟊$\bigstar$ 0.33 0.23 0.10 $\clubsuit$ 0.24 0.17 0.71 🟊$\bigstar$
ST [49] 0.14 0.11 2.87 🟊$\bigstar$ 0.62 0.29 -0.99 $\spadesuit$ 0.44 0.25 -0.20 $\clubsuit$
WideResNet-50 [55] AT [54] 0.09 0.08 4.66 🟊$\bigstar$ 0.50 0.25 -0.94 $\spadesuit$ 0.50 0.25 -0.85 $\spadesuit$
ST [55] 0.06 0.05 6.86 🟊$\bigstar$ 0.83 0.15 -3.92 $\spadesuit$ 0.52 0.18 -0.73 $\spadesuit$
ViT-B [52] AT [54] 0.06 0.04 6.96 🟊$\bigstar$ 0.56 0.21 -1.60 $\spadesuit$ 0.47 0.22 -0.52 $\spadesuit$
ST [52] 0.17 0.12 2.28 🟊$\bigstar$ 0.78 0.25 -1.95 $\spadesuit$ 0.64 0.22 -1.11 $\spadesuit$
TABLE VIII: Performance of Robust (Adversarially Trained: AT) and Non-Robust (Standard Trained: ST) architectures across varied visual patterns - Pattern A (🟊normal-$\bigstar$), Pattern B (normal-$\clubsuit$), Pattern C (normal-$\spadesuit$), and $1,000$ ImageNet1k Classes. The table presents key statistics, including an average of Mean ($\mu_{k^{*}}$), Skewness coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and Number of Classes (N).
Average of Classes with Pattern A (🟊normal-$\bigstar$) (Fractured) Average of Classes with Pattern B (normal-$\clubsuit$) (Overlapped) Average of Classes with Pattern C (normal-$\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
Architecture Type $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc
ResNet-50 [49] AT [54] 0.06 5.77 64.52 975 0.35 0.04 94.60 20 0.57 -1.22 95.60 5 0.07 5.62 65.28
ST [49] 0.10 4.51 74.88 935 0.37 0.14 93.96 56 0.61 -1.04 97.11 9 0.12 4.22 76.15
WideResNet-50 [55] AT [54] 0.08 5.12 66.53 895 0.35 -0.01 89.91 66 0.60 -1.30 95.90 39 0.11 4.53 69.22
ST [55] 0.12 3.85 75.53 650 0.35 -0.02 88.59 135 0.64 -1.70 94.67 215 0.26 2.13 81.41
ViT-B [52] AT [54] 0.08 5.01 68.87 823 0.34 -0.04 88.29 83 0.59 -1.49 94.45 94 0.15 3.98 72.88
ST [52] 0.12 3.77 68.12 652 0.35 0.05 84.97 149 0.67 -1.59 92.73 199 0.26 2.15 75.53
IV-B Analysis of Latent Space of Different Neural Architectures
We have the option to visualize the latent space of ResNet-50 in two ways: using k* distribution, as illustrated in Figure 4 (Left), or employing dimensionality reduction techniques, showcased in Figure 4 (Right). Intriguingly, the t-SNE (Top Left) and UMAP (Bottom Right) visualizations indicate a highly clustered distribution of samples in latent space for ResNet-50, while Isomap (Top Right) and PCA (Bottom Left) show an overlapping distribution of samples in latent space. This creates uncertainty in distinguishing which classes are well-represented and which ones are fragmented in the distribution of samples in latent space.
To address this ambiguity and gain a clearer understanding of the local latent space, we turn to the k* distribution, visualized in Figure 4. By utilizing the k* distribution, we can distinctly identify that the local spaces of the six visualized classes differ. For instance, the Airplane and Elephant classes exhibit more Pattern C (normal-$\spadesuit$) clustered distribution of samples in latent space with a negatively skewed k* distribution.
In contrast, Bicycle and Dog classes showcase Pattern A (🟊normal-$\bigstar$) fractured distribution of samples in latent space with a positively skewed k* distribution, while Car and Truck have Pattern B (normal-$\clubsuit$) overlapped distribution of samples in latent space with an almost uniform k* distribution.
Furthermore, we can perform a comprehensive comparison of various neural architectures, as presented in Table I. This allows us to observe how specific classes are distributed differently across architectures. Notably, the Bicycle class displays Pattern A (🟊normal-$\bigstar$) (fractured distribution of samples) in ResNet and ResNeXt architectures, Pattern B (normal-$\clubsuit$)  (overlapped distribution of samples) in ViT, and Pattern C (normal-$\spadesuit$) (clustered distribution of samples) in EfficientNet-B0. Similarly, the Truck class is Pattern B (normal-$\clubsuit$) (overlapped distribution of samples) in ResNet and ResNeXt architectures, and Pattern C (normal-$\spadesuit$) (clustered distribution of samples) in ViT and EfficientNet-B0. Conversely, Airplane, Dog, and Elephant classes exhibit consistent distribution across all architectures.
Additionally, a comprehensive comparison of neural architectures can be made by calculating the averages of Fractured, Overlapped, and Clustered classes, as detailed in Table II. The results indicate distinct distributions of classes across various neural architectures. For instance, ResNet tends to fracture the latent space, with only 9 clustered classes. In contrast, ViT leans towards clustering of samples in the latent space, with a significant number of classes ($243$) exhibiting clustering.
A general trend is also observed, that there is an increase in the mean of k* distribution $\mu_{k^{*}}$ with improvements in model accuracy. Additionally, a declining trend in the skewness coefficient of k* distribution $\gamma_{k^{*}}$ suggests a transition from fractured to more overlapped distribution of samples in latent space across models.
IV-C Analysis of Latent Space of Different Layers of a Network
To gain insights into the representation of classes across different layers of the same model, we conduct an analysis of the latent space within various layers of ResNet-50. Specifically, we examine the latent space of the final logit layer, the average pooling layer, and several convolution layers situated at the conclusion of different stages in the ResNet-50 architecture, as depicted in Tables III and IV.
Upon observation, we note that the final logit layer induces changes in the representation of certain overlapped classes, transforming them from an average pooling state to predominantly clustered and often fractured representations. Additionally, we note that layers proximal to the input exhibit a highly fractured distribution of samples in latent space. This suggests that layers closer to the input aim to identify abstract patterns in the data, irrespective of the sample’s label as also suggested by Hohman et al. [24]. We also note the association of samples to the supervised class label increases through the network with the best association occurring at the output layer of the network. In contrast, the final layers of neural networks primarily function to cluster and associate these abstract patterns with the sample’s label.
IV-D Analysis of Latent Space of Different Training Distributions
It is well-established that the learned latent space of a neural network undergoes changes based on the distribution of the training data. In order to assess these changes in the learned latent space with respect to training data distribution, we conduct an evaluation using ResNet-50 trained on different data distributions.
Specifically, we compare the models trained on standard ImageNet1k samples, those trained on the standard ImageNet1k dataset [45], a stylized version of ImageNet-1k (Stylized ImageNet) [53], and a combination of both, as illustrated in Tables V and VI. Through this comparison, we can discern alterations in the representation of different classes in the latent space. Notably, training exclusively with Stylized ImageNet results in a more fractured distribution of samples in latent space for the non-stylized images.
Furthermore, we observe a consistent trend where models exhibit better accuracy with higher $\mu_{k^{*}}$ and lower $\gamma_{k^{*}}$ mirroring the pattern observed in Table II. This suggests a correlation between improved model performance and specific characteristics of the k* distributions.
IV-E Analysis of Latent Space of Adversarially Robust Models
TABLE IX: Multi-category evaluation on samples transformed with image crop by using statistical metrics across object categories. Metrics include Mean ($\mu_{k^{*}}$), Standard deviation ($\sigma_{k^{*}}$), Skewness Coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and prevailing Pattern (Pat) in the context of diverse visual patterns.
Airplane Bicycle Car
Image Size ($s$) After Crop $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat
20 0.02 0.01 8.70 5.00 🟊$\bigstar$ 0.01 0.01 8.85 0.00 🟊$\bigstar$ 0.01 0.01 8.83 0.00 🟊$\bigstar$
60 0.06 0.06 6.05 80.00 🟊$\bigstar$ 0.02 0.02 8.56 48.75 🟊$\bigstar$ 0.04 0.05 7.03 36.25 🟊$\bigstar$
100 0.15 0.11 2.54 95.00 🟊$\bigstar$ 0.05 0.06 6.29 77.50 🟊$\bigstar$ 0.13 0.14 2.14 72.50 🟊$\bigstar$
140 0.36 0.22 -0.05 97.50 $\clubsuit$ 0.10 0.09 3.93 92.50 🟊$\bigstar$ 0.28 0.23 0.59 83.75 🟊$\bigstar$
180 0.47 0.24 -0.48 98.75 $\clubsuit$ 0.12 0.09 3.82 96.25 🟊$\bigstar$ 0.40 0.26 0.05 95.00 $\clubsuit$
220 0.62 0.27 -0.98 100.00 $\spadesuit$ 0.16 0.12 2.50 96.25 🟊$\bigstar$ 0.47 0.26 -0.25 93.75 $\clubsuit$
Dog Elephant Truck
Image Size ($s$) After Crop $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat
20 0.02 0.01 8.85 18.75 🟊$\bigstar$ 0.02 0.01 8.78 16.25 🟊$\bigstar$ 0.01 0.01 8.85 7.50 🟊$\bigstar$
60 0.02 0.01 8.70 66.25 🟊$\bigstar$ 0.05 0.05 6.43 81.25 🟊$\bigstar$ 0.03 0.02 8.39 47.50 🟊$\bigstar$
100 0.04 0.03 7.81 82.50 🟊$\bigstar$ 0.20 0.16 1.23 88.75 🟊$\bigstar$ 0.08 0.09 3.87 78.75 🟊$\bigstar$
140 0.08 0.07 5.20 90.00 🟊$\bigstar$ 0.43 0.27 -0.16 97.50 $\clubsuit$ 0.18 0.18 1.40 92.50 🟊$\bigstar$
180 0.13 0.10 2.99 93.75 🟊$\bigstar$ 0.57 0.28 -0.56 98.75 $\spadesuit$ 0.39 0.23 -0.09 97.50 $\clubsuit$
220 0.14 0.11 2.88 96.25 🟊$\bigstar$ 0.61 0.29 -0.91 98.75 $\spadesuit$ 0.45 0.24 -0.21 98.75 $\clubsuit$
TABLE X: Performance on samples transformed with image crop across varied visual patterns - Pattern A (🟊normal-$\bigstar$), Pattern B (normal-$\clubsuit$), Pattern C (normal-$\spadesuit$), and $1,000$ ImageNet1k Classes. The table presents key statistics, including an average of Mean ($\mu_{k^{*}}$), Skewness coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and Number of Classes (N).
Average of Classes with Pattern A (🟊normal-$\bigstar$) (Fractured) Average of Classes with Pattern B (normal-$\clubsuit$) (Overlapped) Average of Classes with Pattern C (normal-$\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
Image Size ($s$) After Crop $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc
20 0.02 6.99 1.28 1,000 0 0 0.02 6.99 1.28
60 0.02 6.91 15.84 1,000 0 0 0.02 6.91 15.84
100 0.03 6.58 38.64 1,000 0 0 0.03 6.58 38.64
140 0.05 6.05 54.24 997 0.38 0.09 94.00 3 0 0.05 6.03 54.36
180 0.06 5.53 62.29 992 0.34 0.12 89.43 7 0.60 -0.85 98.00 1 0.06 5.48 62.52
220 0.07 5.13 67.37 985 0.36 0.11 93.45 11 0.55 -0.76 94.50 4 0.08 5.05 67.77
Having explored the changes in representation space with variations in training distribution, it is essential to address the susceptibility of models to adversarial attacks. To enhance robustness against such attacks, adversarially trained models have been proposed, incorporating adversarial samples in the training distribution [48]. In order to evaluate the shifts in the learned latent space between adversarially trained models and their non-robust counterparts, we examine the latent space of different robust models alongside their non-robust counterparts, as presented in Tables VII and VIII.
Observations from the table indicate that adversarially trained models tend to exhibit a more fractured distribution of samples in latent space compared to their non-robust counterparts. This explains the current tradeoff that exists between accuracy and robustness in the image classification models studied by Tsipras et al. [56]. This suggests that, in an effort to achieve clustering of heterogeneous feature samples within a single cluster as supervised with the pre-defined class label, models often compromise on learning robust features. Additionally, we observe a consistent trend where accuracy is associated with higher $\mu_{k^{*}}$ and lower $\gamma_{k^{*}}$, explaining why adversarially trained models may demonstrate reduced performance on clean samples compared to their non-robust counterparts.
IV-F Analysis of Latent Space of Different Input Transformations
TABLE XI: Multi-category evaluation on samples transformed with image rotation by using statistical metrics across object categories. Metrics include Mean ($\mu_{k^{*}}$), Standard deviation ($\sigma_{k^{*}}$), Skewness Coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and prevailing Pattern (Pat) in the context of diverse visual patterns.
Airplane Bicycle Car
[…]
30 0.10 0.10 3.54 48.75 🟊$\bigstar$ 0.06 0.06 5.74 71.25 🟊$\bigstar$ 0.10 0.11 3.34 63.75 🟊$\bigstar$
60 0.06 0.05 6.39 27.50 🟊$\bigstar$ 0.04 0.04 7.65 42.50 🟊$\bigstar$ 0.05 0.05 6.55 13.75 🟊$\bigstar$
90 0.10 0.10 3.60 50.00 🟊$\bigstar$ 0.08 0.08 4.53 85.00 🟊$\bigstar$ 0.07 0.06 5.64 28.75 🟊$\bigstar$
120 0.05 0.05 6.74 7.50 🟊$\bigstar$ 0.03 0.02 8.43 21.25 🟊$\bigstar$ 0.04 0.04 7.23 11.25 🟊$\bigstar$
150 0.05 0.04 7.23 12.50 🟊$\bigstar$ 0.02 0.02 8.55 20.00 🟊$\bigstar$ 0.04 0.05 6.88 15.00 🟊$\bigstar$
180 0.14 0.12 2.40 90.00 🟊$\bigstar$ 0.08 0.07 5.07 86.25 🟊$\bigstar$ 0.08 0.07 5.23 68.75 🟊$\bigstar$
210 0.05 0.04 7.41 11.25 🟊$\bigstar$ 0.03 0.02 8.38 21.25 🟊$\bigstar$ 0.04 0.06 6.32 17.50 🟊$\bigstar$
240 0.04 0.03 8.02 13.75 🟊$\bigstar$ 0.03 0.03 7.93 33.75 🟊$\bigstar$ 0.04 0.04 7.35 8.75 🟊$\bigstar$
270 0.11 0.08 4.29 51.25 🟊$\bigstar$ 0.07 0.07 5.50 85.00 🟊$\bigstar$ 0.07 0.07 5.18 26.25 🟊$\bigstar$
300 0.05 0.04 6.96 15.00 🟊$\bigstar$ 0.03 0.02 8.24 42.50 🟊$\bigstar$ 0.06 0.07 5.37 22.50 🟊$\bigstar$
330 0.09 0.08 4.37 45.00 🟊$\bigstar$ 0.08 0.07 5.00 63.75 🟊$\bigstar$ 0.12 0.11 2.85 58.75 🟊$\bigstar$
Dog Elephant Truck
[…]
30 0.09 0.09 4.30 65.00 🟊$\bigstar$ 0.14 0.14 2.03 46.25 🟊$\bigstar$ 0.09 0.09 3.86 67.50 🟊$\bigstar$
60 0.04 0.03 7.64 48.75 🟊$\bigstar$ 0.06 0.07 5.69 25.00 🟊$\bigstar$ 0.04 0.03 7.98 16.25 🟊$\bigstar$
90 0.05 0.05 6.39 77.50 🟊$\bigstar$ 0.12 0.12 2.88 65.00 🟊$\bigstar$ 0.05 0.04 7.14 57.50 🟊$\bigstar$
120 0.03 0.03 7.81 22.50 🟊$\bigstar$ 0.03 0.03 8.02 3.75 🟊$\bigstar$ 0.03 0.02 8.23 8.75 🟊$\bigstar$
150 0.03 0.03 8.17 25.00 🟊$\bigstar$ 0.04 0.03 7.82 3.75 🟊$\bigstar$ 0.04 0.03 8.07 7.50 🟊$\bigstar$
180 0.04 0.04 7.36 70.00 🟊$\bigstar$ 0.13 0.12 2.60 68.75 🟊$\bigstar$ 0.05 0.05 6.83 60.00 🟊$\bigstar$
210 0.03 0.04 7.65 21.25 🟊$\bigstar$ 0.04 0.04 7.38 3.75 🟊$\bigstar$ 0.04 0.03 7.81 8.75 🟊$\bigstar$
240 0.03 0.04 7.55 23.75 🟊$\bigstar$ 0.05 0.06 6.25 3.75 🟊$\bigstar$ 0.04 0.04 7.54 7.50 🟊$\bigstar$
270 0.04 0.04 7.48 77.50 🟊$\bigstar$ 0.12 0.12 2.54 61.25 🟊$\bigstar$ 0.05 0.04 7.46 56.25 🟊$\bigstar$
300 0.05 0.06 5.93 45.00 🟊$\bigstar$ 0.06 0.06 6.03 17.50 🟊$\bigstar$ 0.04 0.03 8.14 33.75 🟊$\bigstar$
330 0.08 0.07 5.08 68.75 🟊$\bigstar$ 0.14 0.14 2.07 47.50 🟊$\bigstar$ 0.07 0.08 4.72 57.50 🟊$\bigstar$
TABLE XII: Performance on samples transformed with image rotation across varied visual patterns - Pattern A (🟊normal-$\bigstar$), Pattern B (normal-$\clubsuit$), Pattern C (normal-$\spadesuit$), and $1,000$ ImageNet1k Classes. The table presents key statistics, including an average of Mean ($\mu_{k^{*}}$), Skewness coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and Number of Classes (N).
Average of Classes with Pattern A (🟊normal-$\bigstar$) (Fractured) Average of Classes with Pattern B (normal-$\clubsuit$) (Overlapped) Average of Classes with Pattern C (normal-$\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
[…]
30 0.06 5.51 53.26 983 0.35 0.22 86.53 15 0.58 -0.96 92.00 2 0.07 5.42 53.84
60 0.05 6.05 40.57 993 0.38 0.05 88.33 6 0.50 -0.93 86.00 1 0.05 6.00 40.90
90 0.05 5.94 50.46 988 0.36 0.24 92.73 11 0.54 -1.15 90.00 1 0.06 5.87 50.97
120 0.04 6.38 28.87 999 0.42 -0.41 88.00 1 0 0.04 6.37 28.93
150 0.04 6.39 29.07 997 0.33 0.30 85.00 2 0.48 -0.63 90.00 1 0.04 6.37 29.24
180 0.05 5.95 51.19 991 0.37 0.18 92.25 8 0.53 -0.77 90.00 1 0.05 5.90 51.55
210 0.04 6.38 28.61 999 0.43 -0.46 88.00 1 0 0.04 6.38 28.67
240 0.04 6.37 29.24 999 0.43 -0.40 88.00 1 0 0.04 6.36 29.30
270 0.05 5.95 49.92 989 0.35 0.22 90.80 10 0.51 -0.92 92.00 1 0.06 5.88 50.37
300 0.05 6.05 39.58 994 0.35 0.28 88.40 5 0.49 -0.70 88.00 1 0.05 6.01 39.87
330 0.06 5.51 53.39 978 0.36 0.14 87.68 19 0.58 -0.81 96.00 3 0.07 5.39 54.17
TABLE XIII: Multi-category evaluation on samples transformed with Gaussian noise by using statistical metrics across object categories. Metrics include Mean ($\mu_{k^{*}}$), Standard deviation ($\sigma_{k^{*}}$), Skewness Coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and prevailing Pattern (Pat) in the context of diverse visual patterns.
Airplane Bicycle Car
Strength ($\alpha$) of Gaussian Noise $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat
0.05 0.39 0.23 -0.10 100.00 $\clubsuit$ 0.14 0.11 3.01 93.75 🟊$\bigstar$ 0.38 0.25 0.07 87.50 $\clubsuit$
0.06 0.32 0.20 0.20 100.00 $\clubsuit$ 0.13 0.10 3.21 93.75 🟊$\bigstar$ 0.36 0.24 0.21 87.50 $\clubsuit$
0.08 0.34 0.22 0.08 97.50 $\clubsuit$ 0.12 0.09 3.57 96.25 🟊$\bigstar$ 0.34 0.24 0.36 83.75 $\clubsuit$
0.12 0.20 0.16 1.46 95.00 🟊$\bigstar$ 0.10 0.07 4.92 91.25 🟊$\bigstar$ 0.25 0.21 0.86 85.00 🟊$\bigstar$
0.18 0.13 0.12 2.60 87.50 🟊$\bigstar$ 0.12 0.12 2.77 83.75 🟊$\bigstar$ 0.24 0.20 0.77 77.50 🟊$\bigstar$
0.26 0.07 0.07 5.32 63.75 🟊$\bigstar$ 0.07 0.09 4.11 60.00 🟊$\bigstar$ 0.13 0.12 2.55 61.25 🟊$\bigstar$
0.38 0.03 0.04 7.65 15.00 🟊$\bigstar$ 0.02 0.02 8.56 20.00 🟊$\bigstar$ 0.07 0.08 4.67 27.50 🟊$\bigstar$
0.70 0.02 0.02 8.59 0.00 🟊$\bigstar$ 0.01 0.00 8.86 1.25 🟊$\bigstar$ 0.02 0.01 8.67 2.50 🟊$\bigstar$
1.00 0.02 0.01 8.77 0.00 🟊$\bigstar$ 0.02 0.00 8.86 0.00 🟊$\bigstar$ 0.02 0.01 8.84 0.00 🟊$\bigstar$
Dog Elephant Truck
Strength ($\alpha$) of Gaussian Noise $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat
0.05 0.09 0.08 4.51 91.25 🟊$\bigstar$ 0.51 0.29 -0.37 97.50 $\clubsuit$ 0.32 0.22 0.35 96.25 $\clubsuit$
0.06 0.08 0.08 4.77 87.50 🟊$\bigstar$ 0.42 0.28 0.03 97.50 $\clubsuit$ 0.32 0.22 0.28 97.50 $\clubsuit$
0.08 0.06 0.05 6.25 88.75 🟊$\bigstar$ 0.39 0.27 0.05 96.25 $\clubsuit$ 0.31 0.23 0.47 97.50 $\clubsuit$
0.12 0.04 0.04 7.41 81.25 🟊$\bigstar$ 0.22 0.20 0.97 93.75 🟊$\bigstar$ 0.22 0.19 1.01 97.50 🟊$\bigstar$
0.18 0.03 0.03 8.19 68.75 🟊$\bigstar$ 0.15 0.15 2.03 92.50 🟊$\bigstar$ 0.11 0.11 2.96 93.75 🟊$\bigstar$
0.26 0.02 0.02 8.64 42.50 🟊$\bigstar$ 0.06 0.08 5.07 77.50 🟊$\bigstar$ 0.07 0.08 4.71 67.50 🟊$\bigstar$
0.38 0.02 0.01 8.80 20.00 🟊$\bigstar$ 0.02 0.02 8.43 48.75 🟊$\bigstar$ 0.02 0.02 8.57 21.25 🟊$\bigstar$
0.70 0.02 0.01 8.83 0.00 🟊$\bigstar$ 0.01 0.01 8.86 1.25 🟊$\bigstar$ 0.02 0.01 8.72 0.00 🟊$\bigstar$
1.00 0.01 0.01 8.86 0.00 🟊$\bigstar$ 0.01 0.00 8.86 0.00 🟊$\bigstar$ 0.02 0.01 8.75 0.00 🟊$\bigstar$
TABLE XIV: Performance on samples transformed with Gaussian noise across varied visual patterns - Pattern A (🟊normal-$\bigstar$), Pattern B (normal-$\clubsuit$), Pattern C (normal-$\spadesuit$), and $1,000$ ImageNet1k Classes. The table presents key statistics, including an average of Mean ($\mu_{k^{*}}$), Skewness coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and Number of Classes (N).
Average of Classes with Pattern A (🟊normal-$\bigstar$) (Fractured) Average of Classes with Pattern B (normal-$\clubsuit$) (Overlapped) Average of Classes with Pattern C (normal-$\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
Strength ($\alpha$) of Gaussian Noise $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc
0.05 0.09 4.75 72.26 956 0.35 0.20 93.06 36 0.58 -0.83 96.25 8 0.10 4.55 73.20
0.06 0.09 4.84 71.57 961 0.35 0.18 92.73 33 0.58 -0.83 96.33 6 0.10 4.66 72.41
0.08 0.08 4.96 69.81 972 0.37 0.12 93.00 24 0.59 -0.84 95.50 4 0.09 4.82 70.47
0.12 0.07 5.27 65.31 984 0.38 0.06 93.43 14 0.60 -0.96 93.00 2 0.08 5.19 65.76
0.18 0.05 5.78 56.24 991 0.35 0.20 88.25 8 0.49 -0.74 88.00 1 0.06 5.73 56.53
0.26 0.04 6.30 42.35 998 0.42 -0.11 89.00 2 0 0.04 6.28 42.44
0.38 0.03 6.73 24.49 1,000 0 0 0.03 6.73 24.49
0.70 0.02 6.96 4.78 1,000 0 0 0.02 6.96 4.78
1.00 0.02 6.98 1.45 1,000 0 0 0.02 6.98 1.45
TABLE XV: Multi-category evaluation of samples transformed with image stylization on models trained with different training distributions by using statistical metrics across object categories. Metrics include Mean ($\mu_{k^{*}}$), Standard deviation ($\sigma_{k^{*}}$), Skewness Coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and prevailing Pattern (Pat) in the context of diverse visual patterns.
Airplane Bicycle Car
Training Distribution $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat
ImageNet1k [45] 0.02 0.01 6.94 18.00 🟊$\bigstar$ 0.03 0.02 6.84 46.00 🟊$\bigstar$ 0.04 0.03 6.42 30.00 🟊$\bigstar$
Stylised ImageNet [53] 0.55 0.27 -0.40 94.00 $\clubsuit$ 0.33 0.20 0.40 94.00 $\clubsuit$ 0.18 0.15 1.85 88.00 🟊$\bigstar$
ImageNet1k [45] + Stylised [53] 0.51 0.26 -0.33 94.00 $\clubsuit$ 0.35 0.22 0.32 96.00 $\clubsuit$ 0.15 0.12 2.54 96.00 🟊$\bigstar$
Dog Elephant Truck
Training Distribution $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat
ImageNet1k [45] 0.02 0.01 6.95 34.00 🟊$\bigstar$ 0.02 0.01 6.98 46.00 🟊$\bigstar$ 0.02 0.01 6.91 32.00 🟊$\bigstar$
Stylised [53] 0.16 0.11 2.85 94.00 🟊$\bigstar$ 0.44 0.24 0.21 98.00 $\clubsuit$ 0.09 0.09 4.16 84.00 🟊$\bigstar$
ImageNet1k [45] + Stylised [53] 0.14 0.09 3.69 96.00 🟊$\bigstar$ 0.44 0.21 -0.26 96.00 $\clubsuit$ 0.11 0.09 3.68 82.00 🟊$\bigstar$
TABLE XVI: Performance by samples transformed with image stylization on models trained with different training distributions across varied visual patterns - Pattern A (🟊normal-$\bigstar$), Pattern B (normal-$\clubsuit$), Pattern C (normal-$\spadesuit$), and $1,000$ ImageNet1k Classes. The table presents key statistics, including an average of Mean ($\mu_{k^{*}}$), Skewness coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and Number of Classes (N).
Average of Classes with Pattern A (🟊normal-$\bigstar$) (Fractured) Average of Classes with Pattern B (normal-$\clubsuit$) (Overlapped) Average of Classes with Pattern C (normal-$\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
Training Distribution $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc
ImageNet1k [45] 0.02 6.91 20.18 1,000 0 0 0.02 6.91 20.18
Stylised [53] 0.04 6.25 54.04 999 0.29 0.45 86.00 1 0 0.04 6.24 54.07
ImageNet1k [45] + Stylised [53] 0.04 6.41 47.97 1,000 0 0 0.04 6.41 47.97
TABLE XVII: Multi-category evaluation of adversarial samples created with PGD Attack by using statistical metrics across object categories. Metrics include Mean ($\mu_{k^{*}}$), Standard deviation ($\sigma_{k^{*}}$), Skewness Coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and prevailing Pattern (Pat) in the context of diverse visual patterns.
Airplane Bicycle Car
Architecture
$\epsilon$
$\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat
ResNet-50 Adversarially Trained [54] 2/255 0.33 0.22 -0.08 $\clubsuit$ 0.10 0.08 4.29 🟊$\bigstar$ 0.26 0.19 0.74 🟊$\bigstar$
4/255 0.31 0.22 0.03 $\clubsuit$ 0.09 0.08 4.22 🟊$\bigstar$ 0.17 0.13 2.00 🟊$\bigstar$
8/255 0.26 0.20 0.34 $\clubsuit$ 0.06 0.06 5.65 🟊$\bigstar$ 0.10 0.07 4.97 🟊$\bigstar$
ResNet-50 Standard Trained [49] 2/255 0.04 0.03 7.69 🟊$\bigstar$ 0.02 0.02 8.61 🟊$\bigstar$ 0.18 0.15 1.22 🟊$\bigstar$
4/255 0.06 0.06 5.83 🟊$\bigstar$ 0.02 0.02 8.60 🟊$\bigstar$ 0.29 0.22 0.10 $\clubsuit$
8/255 0.05 0.04 6.94 🟊$\bigstar$ 0.03 0.02 8.37 🟊$\bigstar$ 0.36 0.27 -0.30 $\clubsuit$
Dog Elephant Truck
Architecture
$\epsilon$
$\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat
ResNet-50 Adversarially Trained [54] 2/255 0.04 0.04 7.43 🟊$\bigstar$ 0.19 0.17 1.23 🟊$\bigstar$ 0.17 0.14 1.56 🟊$\bigstar$
4/255 0.03 0.03 7.96 🟊$\bigstar$ 0.10 0.10 3.29 🟊$\bigstar$ 0.12 0.12 2.51 🟊$\bigstar$
8/255 0.02 0.02 8.57 🟊$\bigstar$ 0.06 0.05 6.26 🟊$\bigstar$ 0.07 0.08 4.80 🟊$\bigstar$
ResNet-50 Standard Trained [49] 2/255 0.02 0.01 8.79 🟊$\bigstar$ 0.04 0.04 7.38 🟊$\bigstar$ 0.03 0.02 8.35 🟊$\bigstar$
4/255 0.02 0.02 8.60 🟊$\bigstar$ 0.05 0.05 6.37 🟊$\bigstar$ 0.04 0.04 7.54 🟊$\bigstar$
8/255 0.04 0.04 7.42 🟊$\bigstar$ 0.05 0.05 6.61 🟊$\bigstar$ 0.04 0.04 7.53 🟊$\bigstar$
TABLE XVIII: Performance on adversarial samples created with PGD Attack across varied visual patterns - Pattern A (🟊normal-$\bigstar$), Pattern B (normal-$\clubsuit$), Pattern C (normal-$\spadesuit$), and $1,000$ ImageNet1k Classes. The table presents key statistics, including an average of Mean ($\mu_{k^{*}}$), Skewness coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and Number of Classes (N).
Average of Classes with Pattern A (🟊normal-$\bigstar$) (Fractured) Average of Classes with Pattern B (normal-$\clubsuit$) (Overlapped) Average of Classes with Pattern C (normal-$\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
Architecture
$\epsilon$
$\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc
ResNet-50 Adversarially Trained [54] 2/255 0.05 6.14 50.61 983 0.34 0.05 90.43 14 0.57 -1.18 91.33 3 0.05 6.03 51.29
4/255 0.04 6.38 36.18 990 0.32 0.14 85.50 8 0.57 -1.17 88.00 2 0.04 6.31 36.68
8/255 0.03 6.64 16.45 996 0.30 0.24 72.67 3 0.50 -1.00 90.00 1 0.03 6.62 16.69
ResNet-50 Standard Trained [49] 2/255 0.05 5.82 0.00 980 0.32 0.01 0.00 15 0.50 -0.80 0.00 5 0.06 5.70 0.00
4/255 0.05 5.93 0.00 986 0.33 -0.01 0.00 11 0.50 -0.70 0.00 3 0.05 5.85 0.00
8/255 0.04 6.38 0.03 996 0.31 0.25 0.00 4 0 0.04 6.36 0.03
The learned latent space of neural networks is known to be highly susceptible to slight perturbations in input samples, making it brittle. To comprehend the distribution of such transformed samples within the learned space, we analyze the latent space of ResNet-50 using samples modified with various input transformations. Specifically, we apply five different transformations, described below, to the input sample and individually evaluate their impact. Each sample is transformed once using the transformation to estimate their distribution in the latent space.
Image Crop:
In this transformation, the input sample is center-cropped to a size $s$. The reduced image is then resized to the standard $256\times 256$ dimensions and processed by the neural network. The results of this transformation are presented in Tables IX and X. Notably, as more pixels are cropped from the samples, the distribution of transformed samples becomes more fractured in the latent space. This suggests that the model relies on detecting multiple features within the image, and missing information can lead the model to perceive the sample as entirely different from the original.
Image Rotation:
[…]
Results of this transformation are presented in Tables XI and XII. Our examination of the distributions using k* distribution of rotated samples reveals a highly fractured distribution of transformed samples in the latent space implying that rotated samples are interpreted as distinct from non-rotated samples, also evident from the performance, as measured in accuracy.
Gaussian Noise
: In this transformation, Gaussian noise is added to the input sample. Mathematically, if $x$ is the input sample, $z\sim{\mathcal{N}}(\mu,\alpha^{2})$ is sampled Gaussian noise, and $\alpha$ is the strength of Gaussian Noise, transformed image $x^{\prime}$ can be written as, $x^{\prime}=x+(\alpha\times z)$, which is then processed by the neural network, and the results of this transformation are presented in Tables XIII and XIV.
Similar to the other two input transformations, this also induces fracturing of the distribution of transformed samples in the latent space. Additionally, it is noteworthy that as the strength of the Gaussian noise gradually increases, more classes become fractured, suggesting a gradual breakdown in the features identified by the model.
Image Stylization:
In this transformation, the input sample is deprived of its original texture and replaced with a random painting style, as illustrated by [53]. The results of this transformation on models trained on the standard ImageNet1k [45] dataset, a stylized version of ImageNet-1k (Stylized ImageNet) [53], and a combination of both are displayed in Tables XV and XVI.
Similar to other transformations, stylization in the images induces the fracturing of transformed samples. This further underscores the model’s sensitivity to variations in input samples, leading to distinct representations for stylized samples in the latent space.
Adversarial Perturbation:
To measure the distribution of adversarial samples, we perturb the input sample with adversarial perturbations optimized by a PGD attack [48] with varying strength $\epsilon$ of adversarial perturbation. For a given input sample $x$ and a neural network $f$ such that $f(x)$ is the label of sample $x$ predicted by the neural network, adversarial perturbation $\delta$ can be defined as, $f(x)\neq f(x+\delta)$, where, the adversarial perturbation $\delta$ can be further optimized as,
[…]
The results of adversarial samples on robust and non-robust models are presented in Tables XVII and XVIII. Adversarial perturbations, like other transformations, cause fragmentation in the distribution of samples in the latent space. Although models trained with adversarial techniques exhibit increased robustness and improved transferability as demonstrated by Kotyan et al. [57], they are not completely robust, as indicated by a gradual rise in the fragmentation with stronger perturbations.
In summary, these findings highlight the neural network’s sensitivity to variations in input samples, showcasing that a model perceives transformed samples as substantially different from their original counterparts. The results across various input transformations, including adversarial perturbations, indicate that current models struggle to cluster transformed samples together, interpreting them as distinct not only from the original sample but also from each other.
IV-G Effect of Different Distance Metrics on k* Distribution
TABLE XIX: Multi-category evaluation using different distance metrics by using statistical metrics across object categories. Metrics include Mean ($\mu_{k^{*}}$), Standard deviation ($\sigma_{k^{*}}$), Skewness Coefficient ($\gamma_{k^{*}}$), and prevailing Pattern (Pat) in the context of diverse visual patterns.
Airplane Bicycle Car
Distance Metric $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat
Euclidean ($l_{2}$ norm) 0.57 0.25 -0.84 $\spadesuit$ 0.15 0.11 2.60 🟊$\bigstar$ 0.45 0.26 -0.17 $\clubsuit$
CityBlock ($l_{1}$ norm) 0.55 0.25 -0.71 $\spadesuit$ 0.14 0.10 2.99 🟊$\bigstar$ 0.41 0.24 -0.05 $\clubsuit$
Max Norm ($l_{\infty}$ norm) 0.69 0.31 -1.02 $\spadesuit$ 0.31 0.28 0.74 $\clubsuit$ 0.42 0.26 -0.26 $\clubsuit$
Cosine 0.65 0.29 -0.96 $\spadesuit$ 0.23 0.16 1.29 🟊$\bigstar$ 0.45 0.28 -0.05 $\clubsuit$
Dog Elephant Truck
Distance Metric $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat
Euclidean ($l_{2}$ norm) 0.14 0.11 2.88 🟊$\bigstar$ 0.62 0.29 -0.99 $\spadesuit$ 0.44 0.25 -0.20 $\clubsuit$
CityBlock ($l_{1}$ norm) 0.13 0.10 3.12 🟊$\bigstar$ 0.56 0.27 -0.79 $\clubsuit$ 0.39 0.23 0.05 $\clubsuit$
Max Norm ($l_{\infty}$ norm) 0.10 0.07 4.74 🟊$\bigstar$ 0.63 0.28 -1.09 $\spadesuit$ 0.53 0.30 -0.37 $\clubsuit$
Cosine 0.17 0.12 2.11 🟊$\bigstar$ 0.57 0.29 -0.46 $\spadesuit$ 0.52 0.28 -0.43 $\clubsuit$
TABLE XX: Performance using Different Distance Metrics across varied visual patterns - Pattern A (🟊normal-$\bigstar$), Pattern B (normal-$\clubsuit$), Pattern C (normal-$\spadesuit$), and $1,000$ ImageNet1k Classes. The table presents key statistics, including an average of Mean ($\mu_{k^{*}}$), Skewness coefficient ($\gamma_{k^{*}}$), and Number of Classes (N).
Average of Classes with Pattern A (🟊normal-$\bigstar$)  (Fractured) Average of Classes with Pattern B (normal-$\clubsuit$)  (Overlapped) Average of Classes with Pattern C (normal-$\spadesuit$)  (Clustered) Average of 1,000 ImageNet1k Classes
Distance Metric $\mu_{k^{*}}$ $\gamma_{k^{*}}$ N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ N $\mu_{k^{*}}$ $\gamma_{k^{*}}$
Euclidean ($l_{2}$ norm) 0.10 4.51 935 0.37 0.14 56 0.61 -1.04 9 0.12 4.21
CityBlock ($l_{1}$ norm) 0.09 4.59 949 0.37 0.15 42 0.58 -0.88 9 0.11 4.35
Max Norm ($l_{\infty}$ norm) 0.10 4.27 901 0.39 0.10 82 0.63 -1.12 17 0.14 3.83
Cosine 0.10 4.30 915 0.38 0.07 73 0.62 -1.14 12 0.13 3.93
TABLE XXI: Performance using different tasks and architectures across varied visual patterns - Pattern A (🟊normal-$\bigstar$), Pattern B (normal-$\clubsuit$), Pattern C (normal-$\spadesuit$), and Entire Dataset. The table presents key statistics, including an average of Mean ($\mu_{k^{*}}$), Skewness coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), and Number of Classes (N).
Average of Classes with Pattern A (🟊normal-$\bigstar$) (Fractured) Average of Classes with Pattern B (normal-$\clubsuit$) (Overlapped) Average of Classes with Pattern C (normal-$\spadesuit$) (Clustered) Average on Entire Dataset
Task Architecture $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N
Intent Classification on [46] (Text) DeBERTa V3 [58] 0.18 2.68 64.64 24 0.38 -0.05 81.12 14 0.65 -1.29 83.73 21 0.40 0.62 75.34 59
XLM-RoBERTa [58] 0.21 2.04 79.26 22 0.40 0.07 64.00 8 0.69 -1.60 90.04 29 0.47 -0.02 82.49 59
BERT [58] 0.22 2.18 64.81 15 0.40 -0.09 87.29 8 0.71 -1.75 92.38 36 0.54 -0.53 84.68 59
Multilingual-MiniLM [58] 0.20 2.17 80.22 15 0.42 -0.14 82.87 11 0.71 -1.69 91.33 33 0.53 -0.42 86.93 59
Keyword Spotting on [47] (Audio) AST [59] 0.26 1.36 70.00 1 0.43 0.00 58.83 3 0.77 -1.89 95.18 32 0.73 -1.65 91.45 36
Wav2Vec2-Conformer-L [60] 0.22 1.82 92.00 1 0.38 0.01 40.00 2 0.92 -3.81 98.04 33 0.87 -3.44 94.65 36
To evaluate the sensitivity of the nearest neighbor method to different distance metrics and understand their impact on k* distribution and values, we compute the k* distribution using Euclidean ($l_{2}$ norm), City Block ($l_{1}$ norm), Max Norm ($l_{\infty}$ norm), and Cosine Distances. The results from Table XIX reveal the responsiveness of the k* distribution in terms of metrics such as $\mu_{k^{*}}$, $\sigma_{k^{*}}$, and $\gamma_{k^{*}}$. Despite variations, there is substantial agreement on the sample distribution classification across different distance metrics. Having said that, this trend shifts when collectively assessing performance, as shown in Table XX. Here, we observe minimal sensitivity in metric values like $\mu_{k^{*}}$ and $\gamma_{k^{*}}$ to the choice of distance metrics. In other words, the selection of distance metrics does impact the number of classes classified into multiple patterns.
IV-H Going Beyond Image Classification
Exploring beyond the domain of computer vision, we choose to apply k* distribution on models trained for tasks beyond image classification. Specifically, we evaluate models trained for classifying intent from MASSIVE dataset [46] in the domain of natural language processing, and models trained for keyword spotting from the Speech Commands dataset [47] in the domain of speech processing as shown in Table XXI. We note the wider applicability of k* distribution to multiple domains and observe a similar trend of higher $\mu_{k^{*}}$ and lower $\gamma_{k^{*}}$ for better accuracy. Interestingly, we observe that the models trained for intent classification and keyword spotting were less fractured compared to image classification, suggesting that the feature and label association by neural networks for such tasks can be further improved.
V Conclusion
In this article, we introduce the k* distribution, a methodology grounded in the local neighborhood, to assess the distribution of samples in the learned latent space of neural networks. Our experimental findings indicate that the distribution can be primarily categorized into three distinct patterns: Pattern A (🟊normal-$\bigstar$) representing Fractured distribution of samples in latent space: Identified by a positively skewed k* distribution. Pattern B (normal-$\clubsuit$) representing Overlapped distribution of samples in latent space: Characterized by a nearly uniform k* distribution. Pattern C (normal-$\spadesuit$) representing Clustered distribution of samples in latent space: Indicated by a negatively skewed k* distribution. Using k* distribution, we analyzed the learned latent space of different models, with different training data distributions, and of different layers. Further, the learned latent space was tested against various input transformations to understand how the input transformations affect the learned latent space and are associated with the training distribution. We hope this methodology and analysis will help understand the neural networks’ latent spaces and improve the distributions of samples in latent space.
Acknowledgments
This work was supported by JSPS Grant-in-Aid for Challenging Exploratory Research - Grant Number JP22K19814, JST Strategic Basic Research Promotion Program (AIP Accelerated Research) - Grant Number JP22584686, JSPS Research on Academic Transformation Areas (A) - Grant Number JP22H05194.