### v2

# k* Distribution: Evaluating the Latent Space of Deep Neural Networks using Local Neighborhood Analysis

Abstract
Most examinations of neural networks’ learned latent spaces typically employ dimensionality reduction techniques such as t-SNE or UMAP. These methods distort the local neighborhood in the visualization, making it hard to distinguish the structure of a subset of samples in the latent space. In response to this challenge, we introduce the k* distribution and its corresponding visualization technique This method uses local neighborhood analysis to guarantee the preservation of the structure of sample distributions for individual classes within the subset of the learned latent space. This facilitates easy comparison of different k* distributions, enabling analysis of how various classes are processed by the same neural network. Our study reveals three distinct distributions of samples within the learned latent space subset: a) Fractured, b) Overlapped, and c) Clustered, providing a more profound understanding of existing contemporary visualizations. Experiments show that the distribution of samples within the network’s learned latent space significantly varies depending on the class. Furthermore, we illustrate that our analysis can be applied to explore the latent space of diverse neural network architectures, various layers within neural networks, transformations applied to input samples, and the distribution of training and testing data for neural networks. Thus, the k* distribution should aid in visualizing the structure inside neural networks and further foster their understanding.
Index Terms:
Neural Networks, Latent Space Visualization, Local Neighborhood Analysis, Class Representation, Cluster Analysis
I Introduction
Figure 1: Overview of three distinct basic patterns of k* Distribution. Here, we define the k* value of a sample point as the kth-closest neighbor, which differs in class compared to the test point, i.e., the neighbor (sample) which breaks homogeneity in the local neighborhood of the test point. Pattern A (🟊$\bigstar$) which has positively skewed k* distribution (skewed towards low k* value) representing an ‘Fractured’ distribution of samples in latent space; Pattern B ($\clubsuit$) which has almost uniform k* distribution representing a ‘Overlapped’ distribution of samples in latent space; Pattern C ($\spadesuit$) which has negatively skewed k* distribution (skewed towards high k* value) representing a ‘Clustered’ distribution of samples in latent space.
Asignificant portion of neural network research relies on creating tools to comprehend the acquired latent space and unveil the inner workings of neural networks, often viewed as black boxes. Nevertheless, if researchers are equipped with the essential tools to grasp the learned latent space, it becomes feasible to delve deeper, uncovering insights and reasons that can guide further research in neural networks. Analyzing the neural network’s latent space poses challenges due to its intricate non-convex characteristics. However, directly evaluating and comparing the configuration and distribution of data within high-dimensional latent spaces persists despite the development of various visualization tools.
Existing tools often rely on dimensionality reduction techniques like t-SNE [1] and UMAP [2] to create $2$ or $3$-dimensional scatter plots for individual latent spaces, offering a broad overview. They rely on dimensionality reduction methods rooted in manifold learning, which considers specific manifold characteristics and preserves them while modifying other factors and attributes.
Analyzing the configurations and structures of various sample distributions associated with a specific class within the latent space presents a challenge using existing visualization methods. This challenge extends to the comparison of multiple sample distributions, limiting the analysis of the latent spaces. Additionally, visually comparing multiple latent spaces side by side can be overwhelming and confusing, especially when dealing with numerous points and varying degrees of distortion [3, 4, 5, 6, 7].
If there is too much information kept, it is hard to understand. At the same time, if the information is filtered enough, some perspective is always lacking. Therefore, some complementary tools might be crucial in interpreting these complex latent spaces.
Figure 2: Overview of the framework to create k* Distribution. We use the learned features of a neural network to compute k* values of individual evaluated samples and then compute the k* distribution for a particular class.
This article introduces a tool designed to enhance understanding of the distribution of samples associated with ‘classes/features’ in the learned latent space generated by neural networks. The proposed approach involves leveraging the local neighborhood relationships among these samples in the latent space. More precisely, the nearest-neighbor method is applied to the learned latent space to identify a k*-nearest neighbor (sample) within the local neighborhood featuring a different class from the test sample. This process evaluates the disruption of the homogeneity within the local neighborhood of the tested sample. Subsequently, a distribution of k* values referred to as the k* distribution is generated, incorporating all the samples belonging to a given class. Through an analysis of the k* distribution, three distinct patterns of distribution of samples in the learned latent space are identified as illustrated in Figure 1:
Pattern A (🟊$\bigstar$) representing Fractured distribution of samples in latent space,
Pattern B ($\clubsuit$) representing Overlapped distribution of samples in latent space, or
Pattern C ($\spadesuit$) representing Clustered distribution of samples in latent space.
Contributions: This article provides,
[…]
A new interpretation of latent space learned by the neural network, relying on local neighborhood relationships and homogeneity.
[…]
Identification of various distribution patterns of samples in the latent space based on neighborhood characteristics (see Figure 1).
[…]
A model-agnostic latent space analysis of neural networks, focusing on samples from a single class (see Figure 2).
[…]
A method for straightforwardly comparing different classes and understanding how samples from various classes are distributed in the learned latent space (see Figure 3).
II Related Works
[…]
II-A Visualizing the Latent Space of Neural Networks using Dimensionality Reduction Techniques:
Given that neural networks operate in a high-dimensional space, visualizing their latent space directly is challenging. Various researchers employ dimensionality reduction techniques to overcome this limitation to represent the latent space in 2 or 3 dimensions. There are a lot of techniques for visualizing the latent space, like, [1, 2, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Many of these algorithms require hyperparameter tuning for optimal visualization, making it challenging to fairly compare latent spaces of varying dimensions. Additionally, while these visualizations effectively capture some perspective on the structure of the learned latent space, using them to draw a clear comparison between different local structures is challenging [3, 4, 5, 6, 7].
The effectiveness of dimensionality reduction techniques becomes apparent in instances when the latent space is well-organized and has successfully assimilated the intended information. In such cases, these visualizations demonstrate utility by aligning with pre-established interpretations. Conversely, when applied to latent space lacking clearly defined pre-established knowledge, the efficacy diminishes for such visualizations where the lack of known structure leads to a ‘blob of points’ in the visualization [7].
II-B Visualizing Association of Features in the Latent Space of Neural Network:
Various approaches employ visualizations to understand how neural networks interact with features. Analyzing the responses of units in hidden layers to features provides insights into the learned latent space, elucidating which features the network prioritizes [31, 32]. This involves evaluating individual unit responses or combinations of specific inputs [33].
Another visualization method focuses on understanding neural network attention, revealing which parts of an image are emphasized. Salient regions in the input that influence the network’s output are identified using saliency maps [34] and gradient-weighted class-activation maps [35]. These maps quantify conspicuity at each visual field location, guiding the selection of attended locations based on saliency distribution.
Additionally, interpretable concepts, defined as groups of latent variables in the space that are meaningful, may manifest in the latent space, further contributing to our understanding [31, 32, 36, 37, 38]. In natural language processing, these interpretations often provide insight into learning semantic and linguistic concepts provided by the lexical object [39, 40, 41]. These interpretations are often effectively visualized using analytic systems like [42, 43, 44, 45, 46, 47, 48]; however, the challenge to draw a clear comparison between different local structures and multiple latent spaces persists.
We have seen a few methods significantly contribute to the understanding and visualization of latent spaces using neighborhood-based analyses [49, 50, 51, 52], facilitating the development of more sophisticated models and techniques. They have paved the way for metric learning in neural networks, a domain that focuses on learning distance metrics directly from data [53, 54, 55, 56, 57].
We suggest employing the k* distribution to examine the distribution of samples in the neighborhood of a specific sample and analyze the distribution of samples sharing a predefined label. Our proposed approach presents a framework for effectively comparing latent spaces and sub-spaces with diverse sample distributions (see Figure 2). This framework enables meaningful comparisons between the distribution of samples belonging to different classes and multiple latent spaces, offering additional insights into latent spaces beyond what existing visualizations provide. We hope that our proposed visualization technique using k* distribution can further the advances by providing more insightful analyses of the latent space.
III K* Distribution: Analyzing Homogeneity in the Local Neighborhood of Samples
We propose a methodology based on the local neighborhood to analyze the hyper-dimensional latent space learned by neural networks (see Figure 2). The distribution of samples and clusters in the learned latent space is analyzed by associating them with classes. Through this approach, we gain insights into the distribution patterns within the learned latent space and identify the clusters formed (see Figure 1), thereby enhancing our comprehension of the latent space.
In this context, we introduce the k* distribution by exploring the concept of neighborhoods and their utility in validating the local relationship among features in latent space. The k* distribution is constructed by taking a latent space as input and analyzing the neighborhood of a sample (see Figure 3).
The Nearest Neighbor method, a widely recognized non-parametric technique, enables us to understand the positioning of latent variables in space near a latent variable and their corresponding class distribution. This method aids in gauging the relative distances between latent variables and grouping them into clusters. The underlying principle is grounded in the notion that local neighborhoods offer a reasonable estimate of the sample distribution within the latent space.
The index of the kth nearest neighbor that belongs to a different class than the test sample, disrupting the uniformity of the test sample’s neighborhood, is quantified and termed as the k* value. Essentially, a high k* value indicates that the sample is surrounded by similar points that share the same class. In other words, a high k* value indicates that a neighbor of a different class will be situated a considerable distance from the neighborhood of the specified sample. In particular, the homogeneity of each class cluster within the learned latent space can be assessed using k* (see Figure 3). Additionally, one can ascertain whether a cluster remains cohesive for a given class or if it is fragmented across various spatial locations.
The measurement of the nearest neighbor index allows us to address the sparsity inherent in high-dimensional space. Rather than relying on absolute distance values between points, a relative measure, the neighborhood concept, is utilized. This approach facilitates the comparison of two distinct latent spaces with varying dimensionalities. Importantly, the neighborhood concept is dimensionality-independent, making this technique applicable and effective across low and high dimensions.
Mathematically speaking, let us consider a collection of sample-label pairs $X$: $(x_{1},Y_{1}),(x_{2},Y_{2}),...,(x_{n},Y_{n})$ where $x$ is the input sample, and $Y$ is the label for the input sample. Here, the latent space is embedded with such $x$ points. Let $S$ be the set of all samples $x_{p}\in X$ such that they have the same label $c$, i.e.,
$S_{c}=\{x_{i}\mid\forall x_{i}\in X~{}\text{such that}~{}Y_{i}=c\}.$ (1)
The distance $D$ of sample $x_{p}$ from all other samples can be formulated as follows:
$D(x_{p})=\{\text{distance}(x_{p},x_{i})\mid\forall(x_{i},y_{i})\in X\},$ (2)
where $\text{distance}(a,b)$ is the distance between two samples $a$ and $b$. The commonly employed distance function is the Minkowski distance, a generalization of various distance metrics, including Euclidean and City-block distances. The Minkowski distance of order $r$ between two points $a=(a_{1},a_{2}\ldots a_{d})$ and $b=(b_{1},b_{2}\ldots b_{d})$ in $d$ dimensional space is given by,
$\text{distance}(a,b)_{\text{Minkowski},r}=\left(\sum_{i=1}^{d}|a_{i}-b_{i}|^{r}\right)^{1/r},$ (3)
here, it represents City-block distance ($l_{1}$ norm) when $r=1$; it represents Euclidean distance ($l_{2}$ norm) when $r=2$; and it represents Maximum Norm distance ($l_{\infty}$ norm) when $r=\infty$. $k^{\text{th}}$ neighbour sample $x_{k}^{p}$ to sample $x_{p}$ is defined as,
$x_{k}^{p}=x_{q}\quad\text{where,}\quad q\in\operatorname*{arg\,min}_{x_{q}\in P_{i}}~{}\text{distance}(x_{q},x_{p})$ (4)
such $\text{such that}\quad P_{i}=X-\{x_{j}^{p}\mid\forall j<i\}.$
Similarly, we can define a sorted local neighborhood space $N_{p}$ of sample $(x_{p})$ based on the distance,
$N_{p}=(x_{0}^{p},x_{1}^{p}\ldots x_{n}^{p})$ (5)
[…]
$\text{k}^{*}_{p}=\operatorname*{arg\,min}_{(x_{p},Y_{p})}\{x_{i}^{p}\mid x_{i}^{p}\in N_{p},~{}Y_{i}^{p}\neq Y_{p}\},$ (6)
where $i$ is the index of the nearest neighbor, $Y_{p}$ is the label of test sample $x_{p}$ and $Y_{i}^{p}$ is the label of the nearest neighbor (sample) $x_{i}^{p}$ that differs compared to label $Y_{p}$. Then, the k* distribution $\text{k}^{*}(\cdot)$ of class $c$ can be defined as,
$\text{k}^{*}(S_{c})=\left\{\frac{\text{k}^{*}_{p}}{|S_{c}|}\mid\forall x_{p}\in S_{c}\right\},$ (7)
here $|S_{c}|$ is the cardinality of set $S_{c}$ representing the number of samples belonging to class $c$.
Figure 4: Visualization of the distribution of samples in latent space using, (Left) k* distribution, and (Right) Dimensionality Reduction techniques like t-SNE (Top Left), Isomap (Top Right), PCA (Bottom Left), and UMAP (Bottom Right)  of all classes of 16-class-ImageNet for the Logit Layer of ResNet-50 Architecture. Note that the distribution of samples for a particular class is easier to compare using k* distribution than dimensionality reduction techniques.
Based on the defined k* distribution, we can define certain metrics over it as mentioned below,
Mean of k* distribution ($\mu_{k^{*}}$):
$\mu_{k^{*}}=\frac{1}{|S|}\sum\text{k}^{*}(S)$ (8)
Standard Deviation of k* distribution ($\sigma_{k^{*}}$):
$\sigma_{k^{*}}=\left(\frac{1}{|S|}\sum(\text{k}^{*}(S)-\mu_{k^{*}})^{2}\right)^{\frac{1}{2}}$ (9)
Skewness Coefficient of k* distribution ($\gamma_{k^{*}}$):
$\gamma_{k^{*}}=\frac{\frac{1}{|S|}\sum(\text{k}^{*}(S)-\mu_{k^{*}})^{3}}{\left(\frac{1}{|S|}\sum(\text{k}^{*}(S)-\mu_{k^{*}})^{2}\right)^{3/2}}$ (10)
It measures the asymmetry of the k* distribution about its mean $\mu_{k^{*}}$. The skewness coefficient can be positive, negative, or zero. A positive skewness indicates a distribution that is skewed to the left, i.e., towards lower k* metric values, while a negative skewness indicates a distribution that is skewed to the right, i.e., towards higher k* metric values.
Based on the k* distributions, we observe three distinct patterns (Figure 1) of sample distribution in the latent space:
Pattern A (🟊$\bigstar$) Fractured distribution of samples:
[…]
Pattern B ($\clubsuit$) Overlapped distribution of samples:
[…]
Pattern C ($\spadesuit$) Clustered distribution of samples:
[…]
Note: The k* distribution is not a indicator for measuring a neural network’s classification accuracy. The k* distribution visualizes the learned latent space from a local neighborhood perspective, while metrics like accuracy evaluate it from a different perspective. Both multiple well-separated fractured distributions (indication of overclustering) and well-separated homogeneous clusters (indication of optimal clustering) can lead to high classification accuracy.
Limitations: As our approach revolves around identifying surrounding neighbor samples, we inherit the limitations inherent in the nearest-neighbor method. One such trade-off is sacrificing information about the absolute distances between samples and between the distribution of samples.
IV Experimental Results and Visualization
TABLE I: Multi-category evaluation of various neural architectures by using statistical metrics across object categories.
Airplane Bicycle Car
Architectures $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat
ResNet-50 [58] 0.57 0.25 -0.84 100.00 $\spadesuit$ 0.15 0.11 2.60 97.50 🟊$\bigstar$ 0.45 0.26 -0.17 93.75 $\clubsuit$
ResNeXt-101 [59] 0.54 0.23 -0.71 100.00 $\spadesuit$ 0.22 0.17 1.23 100.00 🟊$\bigstar$ 0.50 0.28 -0.32 97.50 $\clubsuit$
EfficientNet-B0 [60] 0.80 0.19 -3.03 100.00 $\spadesuit$ 0.44 0.20 -0.53 96.25 $\spadesuit$ 0.54 0.20 -0.93 96.25 $\spadesuit$
ViT [61] 0.89 0.12 -5.08 100.00 $\spadesuit$ 0.41 0.15 -0.20 98.75 $\clubsuit$ 0.52 0.20 -0.44 93.75 $\clubsuit$
Dog Elephant Truck
Architecture $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat
ResNet-50 [58] 0.14 0.11 2.88 95.00 🟊$\bigstar$ 0.62 0.29 -0.99 98.75 $\spadesuit$ 0.44 0.25 -0.20 98.75 $\clubsuit$
ResNeXt-101 [59] 0.14 0.12 2.52 97.50 🟊$\bigstar$ 0.71 0.27 -1.38 100.00 $\spadesuit$ 0.49 0.24 -0.45 98.75 $\clubsuit$
EfficientNet-B0 [60] 0.05 0.03 7.89 100.00 🟊$\bigstar$ 0.77 0.18 -2.83 100.00 $\spadesuit$ 0.48 0.20 -0.86 98.75 $\spadesuit$
ViT-B [61] 0.05 0.03 8.06 98.75 🟊$\bigstar$ 0.92 0.09 -5.24 100.00 $\spadesuit$ 0.53 0.19 -0.67 100.00 $\spadesuit$
TABLE II: Performance by various neural architectures across varied visual patterns.
Average of Classes with Pattern A (🟊$\bigstar$) (Fractured) Average of Classes with Pattern B ($\clubsuit$) (Overlapped) Average of Classes with Pattern C ($\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
Architecture $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc
ResNet-50 [58] 0.10 4.51 74.88 935 0.37 0.14 93.96 56 0.61 -1.04 97.11 9 0.12 4.21 76.15
ResNeXt-101 [59] 0.11 4.23 75.05 865 0.38 0.08 93.44 104 0.60 -1.12 96.06 31 0.15 3.63 77.62
EfficientNet-B0 [60] 0.10 4.56 74.66 840 0.37 0.01 92.50 105 0.59 -1.40 95.45 55 0.15 3.76 77.68
ViT [61] 0.12 3.83 74.33 604 0.35 0.01 87.78 153 0.64 -1.85 93.59 243 0.28 1.86 81.07
IV-A Common Experimental Setup
Datasets: Initially, we assess the k* distribution using the entire set of $1280$ images from the $16$-class-ImageNet dataset curated by Geirhos et al. [62]. This dataset contains $80$ images per each of the $16$ classes of the dataset. These are the $16$ entry-level categories from MS-COCO that have the highest number of ImageNet classes mapped via the WordNet hierarchy, making them compatible with the $1,000$ classes of ImageNet-1k dataset [63]. The purpose of employing this dataset is to evaluate individual classes organized naturally. In addition to the $16$-class-ImageNet dataset, we evaluate k* distribution comprehensively using $50,000$ validation images of the original ImageNet-1k dataset [63]. This dataset is divided into $1,000$ classes, each with $50$ images.
We apply various transformations to the samples from the $16$-class-ImageNet dataset and ImageNet-1k dataset to further scrutinize the models’ latent spaces. These transformations include Image Cropping, adding Gaussian Noise, rotating the images, and generating Adversarial and Stylized Versions of the samples. Note that the pre-trained models are not trained on these transformations.
Moreover, to assess the k* distribution across various other tasks, we also use various other datasets for different tasks. We use the English subset of the MASSIVE [64] dataset for intent classification. The testing samples contained $2,970$ samples split across $60$ distinct intent classes. For the keyword spotting task, we used the Speech Commands v0.02 dataset [65], which contains $4,890$ samples split across $36$ command categories. We utilize these datasets to analyze (see Figure 2) and comprehend the characteristics associated with each pattern (see Figure 1).
Deep Neural Network Architectures: We analyze the latent space of pre-trained weights from various open-source neural networks referenced as we have used them in the experiments.
Adversarial Attacks: We subject the trained ResNet-50 model to a Projected Gradient Descent (PGD) attack, as outlined in [66], employing a perturbation magnitude of ${\epsilon}=4/255$. This attack aims to gauge the extent of the effective alteration in representation induced by the adversarial samples. Furthermore, we assess the model that has undergone adversarial training using the PGD attack. This evaluation compares and contrasts the robust model’s latent space with its non-robust counterpart.
Metrics: For our analyses of individual classes, we report; Mean ($\mu_{k^{*}}$), Standard deviation ($\sigma_{k^{*}}$), Skewness Coefficient ($\gamma_{k^{*}}$), Accuracy (Acc), Number of Classes (N), and prevailing Pattern (Pat) in the context of diverse visual patterns.
TABLE III: Multi-category evaluation of various layers of ResNet-50 by using statistical metrics across object categories.
Airplane Bicycle Car
Layer Name $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat
Logit Layer 0.57 0.25 -0.84 $\spadesuit$ 0.15 0.11 2.60 🟊$\bigstar$ 0.45 0.26 -0.17 $\clubsuit$
Average Pooling 0.59 0.25 -0.93 $\clubsuit$ 0.15 0.12 2.59 🟊$\bigstar$ 0.42 0.26 -0.07 $\clubsuit$
Stage4 Block3 Conv3 0.27 0.16 0.76 $\clubsuit$ 0.08 0.06 5.50 🟊$\bigstar$ 0.06 0.06 6.12 🟊$\bigstar$
Stage3 Block6 Conv3 0.07 0.08 4.78 🟊$\bigstar$ 0.03 0.02 8.37 🟊$\bigstar$ 0.03 0.03 8.14 🟊$\bigstar$
Stage2 Block4 Conv3 0.01 0.00 8.88 🟊$\bigstar$ 0.01 0.00 8.89 🟊$\bigstar$ 0.01 0.00 8.89 🟊$\bigstar$
Dog Elephant Truck
Layer Name $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat
Logit Layer 0.14 0.11 2.88 🟊$\bigstar$ 0.62 0.29 -0.99 $\spadesuit$ 0.44 0.25 -0.20 $\clubsuit$
Average Pooling 0.09 0.07 4.98 🟊$\bigstar$ 0.62 0.28 -1.06 $\spadesuit$ 0.43 0.25 -0.19 $\clubsuit$
Stage4 Block3 Conv3 0.01 0.00 8.87 🟊$\bigstar$ 0.10 0.08 4.41 🟊$\bigstar$ 0.06 0.05 6.55 🟊$\bigstar$
Stage3 Block6 Conv3 0.02 0.01 8.82 🟊$\bigstar$ 0.03 0.02 8.24 🟊$\bigstar$ 0.02 0.01 8.71 🟊$\bigstar$
Stage2 Block4 Conv3 0.01 0.00 8.89 🟊$\bigstar$ 0.01 0.00 8.89 🟊$\bigstar$ 0.01 0.00 8.89 🟊$\bigstar$
TABLE IV: Performance by different layers of ResNet-50 across varied visual patterns.
Average of Classes with Pattern A (🟊$\bigstar$)  (Fractured) Average of Classes with Pattern B ($\clubsuit$)  (Overlapped) Average of Classes with Pattern C ($\spadesuit$)  (Clustered) Average of 1,000 ImageNet1k Classes
Layer Name $\mu_{k^{*}}$ $\gamma_{k^{*}}$ N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ N $\mu_{k^{*}}$ $\gamma_{k^{*}}$
Logit Layer 0.10 4.51 935 0.37 0.14 56 0.61 -1.04 9 0.12 4.21
Average Pooling 0.09 4.59 837 0.34 0.35 157 0.72 -1.73 6 0.13 3.89
Stage4 Block3 Conv3 0.05 5.90 998 0.33 0.48 2 0 0.05 5.89
Stage3 Block6 Conv3 0.02 6.96 1,000 0 0 0.02 6.96
Stage2 Block4 Conv3 0.02 7.00 1,000 0 0 0.02 7.00
IV-B Analysis of Latent Space of Different Neural Architectures
We have the option to visualize the learned latent space of ResNet-50 (logit layer) in two ways:
a) Using k* distribution, as illustrated in Figure 4 (Left), or,
b) By employing dimensionality reduction techniques, showcased in Figure 4 (Right).
Intriguingly, the t-SNE (Top Left) and UMAP (Bottom Right) visualizations indicate a highly clustered distribution of samples in latent space for ResNet-50, while Isomap (Top Right) and PCA (Bottom Left) show an overlapping distribution of samples in latent space. This creates uncertainty in distinguishing which classes are well-represented and which are fragmented in the distribution of samples in latent space.
To address this ambiguity and better understand the local latent space, we turn to the k* distribution, visualized in Figure 4 (Left). By utilizing the k* distribution, we can distinctly identify that the local spaces of the six visualized classes differ. For instance, the Airplane and Elephant classes exhibit more Pattern C ($\spadesuit$) clustered distribution of samples in latent space with a negatively skewed k* distribution.
In contrast, Bicycle and Dog classes showcase Pattern A (🟊$\bigstar$) fractured distribution of samples in latent space with a positively skewed k* distribution, while Car and Truck have Pattern B ($\clubsuit$) overlapped distribution of samples in latent space with an almost uniform k* distribution.
Furthermore, we can comprehensively compare various neural architectures, as presented in Table I. This allows us to observe how specific classes are distributed differently across architectures. Notably, the Bicycle class displays Pattern A (🟊$\bigstar$) (fractured distribution of samples) in ResNet and ResNeXt architectures, Pattern B ($\clubsuit$)  (overlapped distribution of samples) in ViT, and Pattern C ($\spadesuit$) (clustered distribution of samples) in EfficientNet-B0. Similarly, the Truck class is Pattern B ($\clubsuit$) (overlapped distribution of samples) in ResNet and ResNeXt architectures and Pattern C ($\spadesuit$) (clustered distribution of samples) in ViT and EfficientNet-B0. Conversely, Airplane, Dog, and Elephant classes exhibit consistent distribution across all architectures.
Additionally, a comprehensive comparison of neural architectures can be made by calculating the averages of Fractured, Overlapped, and Clustered classes, as detailed in Table II. The results indicate distinct distributions of classes across various architectures. For instance, ResNet tends to fracture the latent space, with only $9$ clustered classes. In contrast, ViT leans towards clustering of samples in the latent space, with a significant number of classes ($243$) exhibiting clustering, i.e., Pattern A (🟊$\bigstar$).
A general trend is also observed: an increase in the mean of k* distribution $\mu_{k^{*}}$ with improvements in model accuracy. Additionally, a declining general trend in the skewness coefficient of k* distribution $\gamma_{k^{*}}$ suggests a transition from fractured to the more overlapped distribution of samples in latent space across tested models.
IV-C Analysis of Latent Space of Different Layers of a Network
TABLE V: Multi-category evaluation of models trained on different distributions by using statistical metrics across object categories.
Airplane Bicycle Car
Training Distribution $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat
ImageNet1k [63] 0.57 0.25 -0.84 100.00 $\spadesuit$ 0.15 0.11 2.60 97.50 🟊$\bigstar$ 0.45 0.26 -0.17 93.75 $\clubsuit$
Stylised ImageNet [67] 0.41 0.24 -0.29 95.00 $\clubsuit$ 0.13 0.09 3.72 93.75 🟊$\bigstar$ 0.23 0.18 1.12 87.50 🟊$\bigstar$
ImageNet1k + Stylised [67] 0.58 0.24 -0.97 100.00 $\spadesuit$ 0.17 0.13 1.97 98.75 🟊$\bigstar$ 0.43 0.27 -0.02 92.50 $\clubsuit$
Dog Elephant Truck
Training Distribution $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat
ImageNet1k [63] 0.14 0.11 2.88 95.00 🟊$\bigstar$ 0.62 0.29 -0.99 98.75 $\spadesuit$ 0.44 0.25 -0.20 98.75 $\clubsuit$
Stylised ImageNet [67] 0.09 0.07 5.20 95.00 🟊$\bigstar$ 0.41 0.28 0.01 98.75 $\clubsuit$ 0.26 0.21 0.75 97.50 🟊$\bigstar$
ImageNet1k + Stylised [67] 0.16 0.13 2.22 96.25 🟊$\bigstar$ 0.64 0.29 -0.85 100.00 $\spadesuit$ 0.40 0.22 -0.07 100.00 $\clubsuit$
TABLE VI: Performance by models trained on different training datasets across varied visual patterns.
Average of Classes with Pattern A (🟊$\bigstar$) (Fractured) Average of Classes with Pattern B ($\clubsuit$) (Overlapped) Average of Classes with Pattern C ($\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
Training Distribution $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc
ImageNet1k [63] 0.10 4.51 74.88 935 0.37 0.14 93.96 56 0.61 -1.04 97.11 9 0.12 4.21 76.15
Stylised ImageNet [67] 0.05 5.84 59.99 994 0.35 0.19 91.33 6 0 0.06 5.80 60.18
ImageNet1k + Stylised [67] 0.09 4.67 73.36 939 0.37 0.12 92.90 51 0.57 -0.98 96.20 10 0.11 4.38 74.59
TABLE VII: Multi-category evaluation of Robust (Adversarially Trained: AT) and Non-Robust (Standard Trained: ST) models by using statistical metrics across object categories.
Airplane Bicycle Car
Architecture Type $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat
ResNet-50 [58] AT [68] 0.40 0.22 -0.40 $\clubsuit$ 0.13 0.09 3.56 🟊$\bigstar$ 0.35 0.23 0.18 $\clubsuit$
ST [58] 0.57 0.25 -0.84 $\spadesuit$ 0.15 0.11 2.60 🟊$\bigstar$ 0.45 0.26 -0.17 $\clubsuit$
WideResNet-50 [69] AT [68] 0.64 0.26 -1.51 $\spadesuit$ 0.16 0.10 2.76 🟊$\bigstar$ 0.57 0.27 -0.81 $\spadesuit$
ST [69] 0.87 0.17 -3.75 $\spadesuit$ 0.21 0.14 1.30 🟊$\bigstar$ 0.57 0.20 -0.57 $\spadesuit$
ViT-B [61] AT [68] 0.76 0.17 -3.26 $\spadesuit$ 0.25 0.16 0.84 🟊$\bigstar$ 0.62 0.23 -1.26 $\spadesuit$
ST [61] 0.91 0.19 -3.27 $\spadesuit$ 0.30 0.17 0.25 $\clubsuit$ 0.56 0.26 -0.75 $\spadesuit$
Dog Elephant Truck
Architecture Type $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat
ResNet-50 [58] AT [68] 0.06 0.05 6.34 🟊$\bigstar$ 0.33 0.23 0.10 $\clubsuit$ 0.24 0.17 0.71 🟊$\bigstar$
ST [58] 0.14 0.11 2.87 🟊$\bigstar$ 0.62 0.29 -0.99 $\spadesuit$ 0.44 0.25 -0.20 $\clubsuit$
WideResNet-50 [69] AT [68] 0.09 0.08 4.66 🟊$\bigstar$ 0.50 0.25 -0.94 $\spadesuit$ 0.50 0.25 -0.85 $\spadesuit$
ST [69] 0.06 0.05 6.86 🟊$\bigstar$ 0.83 0.15 -3.92 $\spadesuit$ 0.52 0.18 -0.73 $\spadesuit$
ViT-B [61] AT [68] 0.06 0.04 6.96 🟊$\bigstar$ 0.56 0.21 -1.60 $\spadesuit$ 0.47 0.22 -0.52 $\spadesuit$
ST [61] 0.17 0.12 2.28 🟊$\bigstar$ 0.78 0.25 -1.95 $\spadesuit$ 0.64 0.22 -1.11 $\spadesuit$
TABLE VIII: Performance of Robust (Adversarially Trained: AT) and Non-Robust (Standard Trained: ST) models across varied visual patterns.
Average of Classes with Pattern A (🟊$\bigstar$) (Fractured) Average of Classes with Pattern B ($\clubsuit$) (Overlapped) Average of Classes with Pattern C ($\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
Architecture Type $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc
ResNet-50 [58] AT [68] 0.06 5.77 64.52 975 0.35 0.04 94.60 20 0.57 -1.22 95.60 5 0.07 5.62 65.28
ST [58] 0.10 4.51 74.88 935 0.37 0.14 93.96 56 0.61 -1.04 97.11 9 0.12 4.22 76.15
WideResNet-50 [69] AT [68] 0.08 5.12 66.53 895 0.35 -0.01 89.91 66 0.60 -1.30 95.90 39 0.11 4.53 69.22
ST [69] 0.12 3.85 75.53 650 0.35 -0.02 88.59 135 0.64 -1.70 94.67 215 0.26 2.13 81.41
ViT-B [61] AT [68] 0.08 5.01 68.87 823 0.34 -0.04 88.29 83 0.59 -1.49 94.45 94 0.15 3.98 72.88
ST [61] 0.12 3.77 68.12 652 0.35 0.05 84.97 149 0.67 -1.59 92.73 199 0.26 2.15 75.53
We analyze the latent space of various layers within ResNet-50 to gain insights into how classes are represented across different layers of the same model. Specifically, we examine the latent space of the final logit layer, the average pooling layer, and several convolution layers after different stages in the ResNet-50 architecture, as depicted in Tables III and IV.
Upon observation, we note that that number of overlapped classes from the average pooling layer decreases from $153$ to $56$ in the final logit layer, suggesting that logit layer fractures the overlapped regions. This is evident from the results as the number of fractured classes increases from $837$ in average pooling layer to $935$ for the logit layer.
IV-D Analysis of Latent Space of Different Training Distributions
It is well-established that the learned latent space of a neural network undergoes changes based on the distribution of the training data. In order to assess these changes in the learned latent space with respect to training data distribution, we conduct an evaluation using ResNet-50 trained on different data distributions.
Specifically, we compare the models trained on standard ImageNet1k samples, those trained on the standard ImageNet1k dataset [63], a stylized version of ImageNet-1k (Stylized ImageNet) [67], and a combination of both, as illustrated in Tables V and VI. Through this comparison, we can discern alterations in the representation of different classes in the latent space. Notably, training exclusively with Stylized ImageNet results in a more fractured distribution of samples in latent space for the non-stylized images.
Furthermore, we observe a consistent trend where models exhibit better accuracy with higher $\mu_{k^{*}}$ and lower $\gamma_{k^{*}}$ mirroring the pattern observed in Table II. This suggests a correlation between improved model performance and specific characteristics of the k* distributions.
IV-E Analysis of Latent Space of Adversarially Robust Models
TABLE IX: Multi-category evaluation on samples transformed with image crop by using statistical metrics across object categories.
Airplane Bicycle Car
Image Size ($s$) After Crop $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat
20 0.02 0.01 8.70 5.00 🟊$\bigstar$ 0.01 0.01 8.85 0.00 🟊$\bigstar$ 0.01 0.01 8.83 0.00 🟊$\bigstar$
60 0.06 0.06 6.05 80.00 🟊$\bigstar$ 0.02 0.02 8.56 48.75 🟊$\bigstar$ 0.04 0.05 7.03 36.25 🟊$\bigstar$
100 0.15 0.11 2.54 95.00 🟊$\bigstar$ 0.05 0.06 6.29 77.50 🟊$\bigstar$ 0.13 0.14 2.14 72.50 🟊$\bigstar$
140 0.36 0.22 -0.05 97.50 $\clubsuit$ 0.10 0.09 3.93 92.50 🟊$\bigstar$ 0.28 0.23 0.59 83.75 🟊$\bigstar$
180 0.47 0.24 -0.48 98.75 $\clubsuit$ 0.12 0.09 3.82 96.25 🟊$\bigstar$ 0.40 0.26 0.05 95.00 $\clubsuit$
220 0.62 0.27 -0.98 100.00 $\spadesuit$ 0.16 0.12 2.50 96.25 🟊$\bigstar$ 0.47 0.26 -0.25 93.75 $\clubsuit$
Dog Elephant Truck
Image Size ($s$) After Crop $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat
20 0.02 0.01 8.85 18.75 🟊$\bigstar$ 0.02 0.01 8.78 16.25 🟊$\bigstar$ 0.01 0.01 8.85 7.50 🟊$\bigstar$
60 0.02 0.01 8.70 66.25 🟊$\bigstar$ 0.05 0.05 6.43 81.25 🟊$\bigstar$ 0.03 0.02 8.39 47.50 🟊$\bigstar$
100 0.04 0.03 7.81 82.50 🟊$\bigstar$ 0.20 0.16 1.23 88.75 🟊$\bigstar$ 0.08 0.09 3.87 78.75 🟊$\bigstar$
140 0.08 0.07 5.20 90.00 🟊$\bigstar$ 0.43 0.27 -0.16 97.50 $\clubsuit$ 0.18 0.18 1.40 92.50 🟊$\bigstar$
180 0.13 0.10 2.99 93.75 🟊$\bigstar$ 0.57 0.28 -0.56 98.75 $\spadesuit$ 0.39 0.23 -0.09 97.50 $\clubsuit$
220 0.14 0.11 2.88 96.25 🟊$\bigstar$ 0.61 0.29 -0.91 98.75 $\spadesuit$ 0.45 0.24 -0.21 98.75 $\clubsuit$
TABLE X: Performance on samples transformed with image crop across varied visual patterns.
Average of Classes with Pattern A (🟊$\bigstar$) (Fractured) Average of Classes with Pattern B ($\clubsuit$) (Overlapped) Average of Classes with Pattern C ($\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
Image Size ($s$) After Crop $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc
20 0.02 6.99 1.28 1,000 0 0 0.02 6.99 1.28
60 0.02 6.91 15.84 1,000 0 0 0.02 6.91 15.84
100 0.03 6.58 38.64 1,000 0 0 0.03 6.58 38.64
140 0.05 6.05 54.24 997 0.38 0.09 94.00 3 0 0.05 6.03 54.36
180 0.06 5.53 62.29 992 0.34 0.12 89.43 7 0.60 -0.85 98.00 1 0.06 5.48 62.52
220 0.07 5.13 67.37 985 0.36 0.11 93.45 11 0.55 -0.76 94.50 4 0.08 5.05 67.77
TABLE XI: Multi-category evaluation on samples transformed with image rotation by using statistical metrics across object categories.
Airplane Bicycle Car
[…]
30 0.10 0.10 3.54 48.75 🟊$\bigstar$ 0.06 0.06 5.74 71.25 🟊$\bigstar$ 0.10 0.11 3.34 63.75 🟊$\bigstar$
60 0.06 0.05 6.39 27.50 🟊$\bigstar$ 0.04 0.04 7.65 42.50 🟊$\bigstar$ 0.05 0.05 6.55 13.75 🟊$\bigstar$
90 0.10 0.10 3.60 50.00 🟊$\bigstar$ 0.08 0.08 4.53 85.00 🟊$\bigstar$ 0.07 0.06 5.64 28.75 🟊$\bigstar$
120 0.05 0.05 6.74 7.50 🟊$\bigstar$ 0.03 0.02 8.43 21.25 🟊$\bigstar$ 0.04 0.04 7.23 11.25 🟊$\bigstar$
150 0.05 0.04 7.23 12.50 🟊$\bigstar$ 0.02 0.02 8.55 20.00 🟊$\bigstar$ 0.04 0.05 6.88 15.00 🟊$\bigstar$
180 0.14 0.12 2.40 90.00 🟊$\bigstar$ 0.08 0.07 5.07 86.25 🟊$\bigstar$ 0.08 0.07 5.23 68.75 🟊$\bigstar$
210 0.05 0.04 7.41 11.25 🟊$\bigstar$ 0.03 0.02 8.38 21.25 🟊$\bigstar$ 0.04 0.06 6.32 17.50 🟊$\bigstar$
240 0.04 0.03 8.02 13.75 🟊$\bigstar$ 0.03 0.03 7.93 33.75 🟊$\bigstar$ 0.04 0.04 7.35 8.75 🟊$\bigstar$
270 0.11 0.08 4.29 51.25 🟊$\bigstar$ 0.07 0.07 5.50 85.00 🟊$\bigstar$ 0.07 0.07 5.18 26.25 🟊$\bigstar$
300 0.05 0.04 6.96 15.00 🟊$\bigstar$ 0.03 0.02 8.24 42.50 🟊$\bigstar$ 0.06 0.07 5.37 22.50 🟊$\bigstar$
330 0.09 0.08 4.37 45.00 🟊$\bigstar$ 0.08 0.07 5.00 63.75 🟊$\bigstar$ 0.12 0.11 2.85 58.75 🟊$\bigstar$
Dog Elephant Truck
[…]
30 0.09 0.09 4.30 65.00 🟊$\bigstar$ 0.14 0.14 2.03 46.25 🟊$\bigstar$ 0.09 0.09 3.86 67.50 🟊$\bigstar$
60 0.04 0.03 7.64 48.75 🟊$\bigstar$ 0.06 0.07 5.69 25.00 🟊$\bigstar$ 0.04 0.03 7.98 16.25 🟊$\bigstar$
90 0.05 0.05 6.39 77.50 🟊$\bigstar$ 0.12 0.12 2.88 65.00 🟊$\bigstar$ 0.05 0.04 7.14 57.50 🟊$\bigstar$
120 0.03 0.03 7.81 22.50 🟊$\bigstar$ 0.03 0.03 8.02 3.75 🟊$\bigstar$ 0.03 0.02 8.23 8.75 🟊$\bigstar$
150 0.03 0.03 8.17 25.00 🟊$\bigstar$ 0.04 0.03 7.82 3.75 🟊$\bigstar$ 0.04 0.03 8.07 7.50 🟊$\bigstar$
180 0.04 0.04 7.36 70.00 🟊$\bigstar$ 0.13 0.12 2.60 68.75 🟊$\bigstar$ 0.05 0.05 6.83 60.00 🟊$\bigstar$
210 0.03 0.04 7.65 21.25 🟊$\bigstar$ 0.04 0.04 7.38 3.75 🟊$\bigstar$ 0.04 0.03 7.81 8.75 🟊$\bigstar$
240 0.03 0.04 7.55 23.75 🟊$\bigstar$ 0.05 0.06 6.25 3.75 🟊$\bigstar$ 0.04 0.04 7.54 7.50 🟊$\bigstar$
270 0.04 0.04 7.48 77.50 🟊$\bigstar$ 0.12 0.12 2.54 61.25 🟊$\bigstar$ 0.05 0.04 7.46 56.25 🟊$\bigstar$
300 0.05 0.06 5.93 45.00 🟊$\bigstar$ 0.06 0.06 6.03 17.50 🟊$\bigstar$ 0.04 0.03 8.14 33.75 🟊$\bigstar$
330 0.08 0.07 5.08 68.75 🟊$\bigstar$ 0.14 0.14 2.07 47.50 🟊$\bigstar$ 0.07 0.08 4.72 57.50 🟊$\bigstar$
Having explored the changes in representation space with variations in training distribution, it is essential to address models’ susceptibility to adversarial attacks. To enhance robustness against such attacks, adversarially trained models have been proposed, incorporating adversarial samples in the training distribution [66]. In order to evaluate the shifts in the learned latent space between adversarially trained models and their non-robust counterparts, we examine the latent space of different robust models alongside their non-robust counterparts, as presented in Tables VII and VIII.
Observations from the table indicate that adversarially trained models tend to exhibit a more fractured distribution of samples in latent space than their non-robust counterparts. This explains the current trade-off between accuracy and robustness in the image classification models studied by Tsipras et al. [70]. This suggests that, in an effort to achieve clustering of heterogeneous feature samples within a single cluster as supervised with the pre-defined class label, models often compromise on learning robust features. Additionally, we observe a consistent trend where accuracy is associated with higher $\mu_{k^{*}}$ and lower $\gamma_{k^{*}}$, explaining why adversarially trained models may demonstrate reduced performance on clean samples compared to their non-robust counterparts.
IV-F Analysis of Latent Space of Different Input Transformations
TABLE XII: Performance on samples transformed with image rotation across varied visual patterns.
Average of Classes with Pattern A (🟊$\bigstar$) (Fractured) Average of Classes with Pattern B ($\clubsuit$) (Overlapped) Average of Classes with Pattern C ($\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
[…]
30 0.06 5.51 53.26 983 0.35 0.22 86.53 15 0.58 -0.96 92.00 2 0.07 5.42 53.84
60 0.05 6.05 40.57 993 0.38 0.05 88.33 6 0.50 -0.93 86.00 1 0.05 6.00 40.90
90 0.05 5.94 50.46 988 0.36 0.24 92.73 11 0.54 -1.15 90.00 1 0.06 5.87 50.97
120 0.04 6.38 28.87 999 0.42 -0.41 88.00 1 0 0.04 6.37 28.93
150 0.04 6.39 29.07 997 0.33 0.30 85.00 2 0.48 -0.63 90.00 1 0.04 6.37 29.24
180 0.05 5.95 51.19 991 0.37 0.18 92.25 8 0.53 -0.77 90.00 1 0.05 5.90 51.55
210 0.04 6.38 28.61 999 0.43 -0.46 88.00 1 0 0.04 6.38 28.67
240 0.04 6.37 29.24 999 0.43 -0.40 88.00 1 0 0.04 6.36 29.30
270 0.05 5.95 49.92 989 0.35 0.22 90.80 10 0.51 -0.92 92.00 1 0.06 5.88 50.37
300 0.05 6.05 39.58 994 0.35 0.28 88.40 5 0.49 -0.70 88.00 1 0.05 6.01 39.87
330 0.06 5.51 53.39 978 0.36 0.14 87.68 19 0.58 -0.81 96.00 3 0.07 5.39 54.17
TABLE XIII: Multi-category evaluation on samples transformed with Gaussian noise by using statistical metrics across object categories.
Airplane Bicycle Car
Strength ($\alpha$) of Gaussian Noise $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat
0.05 0.39 0.23 -0.10 100.00 $\clubsuit$ 0.14 0.11 3.01 93.75 🟊$\bigstar$ 0.38 0.25 0.07 87.50 $\clubsuit$
0.06 0.32 0.20 0.20 100.00 $\clubsuit$ 0.13 0.10 3.21 93.75 🟊$\bigstar$ 0.36 0.24 0.21 87.50 $\clubsuit$
0.08 0.34 0.22 0.08 97.50 $\clubsuit$ 0.12 0.09 3.57 96.25 🟊$\bigstar$ 0.34 0.24 0.36 83.75 $\clubsuit$
0.12 0.20 0.16 1.46 95.00 🟊$\bigstar$ 0.10 0.07 4.92 91.25 🟊$\bigstar$ 0.25 0.21 0.86 85.00 🟊$\bigstar$
0.18 0.13 0.12 2.60 87.50 🟊$\bigstar$ 0.12 0.12 2.77 83.75 🟊$\bigstar$ 0.24 0.20 0.77 77.50 🟊$\bigstar$
0.26 0.07 0.07 5.32 63.75 🟊$\bigstar$ 0.07 0.09 4.11 60.00 🟊$\bigstar$ 0.13 0.12 2.55 61.25 🟊$\bigstar$
0.38 0.03 0.04 7.65 15.00 🟊$\bigstar$ 0.02 0.02 8.56 20.00 🟊$\bigstar$ 0.07 0.08 4.67 27.50 🟊$\bigstar$
0.70 0.02 0.02 8.59 0.00 🟊$\bigstar$ 0.01 0.00 8.86 1.25 🟊$\bigstar$ 0.02 0.01 8.67 2.50 🟊$\bigstar$
1.00 0.02 0.01 8.77 0.00 🟊$\bigstar$ 0.02 0.00 8.86 0.00 🟊$\bigstar$ 0.02 0.01 8.84 0.00 🟊$\bigstar$
Dog Elephant Truck
Strength ($\alpha$) of Gaussian Noise $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat
0.05 0.09 0.08 4.51 91.25 🟊$\bigstar$ 0.51 0.29 -0.37 97.50 $\clubsuit$ 0.32 0.22 0.35 96.25 $\clubsuit$
0.06 0.08 0.08 4.77 87.50 🟊$\bigstar$ 0.42 0.28 0.03 97.50 $\clubsuit$ 0.32 0.22 0.28 97.50 $\clubsuit$
0.08 0.06 0.05 6.25 88.75 🟊$\bigstar$ 0.39 0.27 0.05 96.25 $\clubsuit$ 0.31 0.23 0.47 97.50 $\clubsuit$
0.12 0.04 0.04 7.41 81.25 🟊$\bigstar$ 0.22 0.20 0.97 93.75 🟊$\bigstar$ 0.22 0.19 1.01 97.50 🟊$\bigstar$
0.18 0.03 0.03 8.19 68.75 🟊$\bigstar$ 0.15 0.15 2.03 92.50 🟊$\bigstar$ 0.11 0.11 2.96 93.75 🟊$\bigstar$
0.26 0.02 0.02 8.64 42.50 🟊$\bigstar$ 0.06 0.08 5.07 77.50 🟊$\bigstar$ 0.07 0.08 4.71 67.50 🟊$\bigstar$
0.38 0.02 0.01 8.80 20.00 🟊$\bigstar$ 0.02 0.02 8.43 48.75 🟊$\bigstar$ 0.02 0.02 8.57 21.25 🟊$\bigstar$
0.70 0.02 0.01 8.83 0.00 🟊$\bigstar$ 0.01 0.01 8.86 1.25 🟊$\bigstar$ 0.02 0.01 8.72 0.00 🟊$\bigstar$
1.00 0.01 0.01 8.86 0.00 🟊$\bigstar$ 0.01 0.00 8.86 0.00 🟊$\bigstar$ 0.02 0.01 8.75 0.00 🟊$\bigstar$
TABLE XIV: Performance on samples transformed with Gaussian noise across varied visual patterns.
Average of Classes with Pattern A (🟊$\bigstar$) (Fractured) Average of Classes with Pattern B ($\clubsuit$) (Overlapped) Average of Classes with Pattern C ($\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
Strength ($\alpha$) of Gaussian Noise $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc
0.05 0.09 4.75 72.26 956 0.35 0.20 93.06 36 0.58 -0.83 96.25 8 0.10 4.55 73.20
0.06 0.09 4.84 71.57 961 0.35 0.18 92.73 33 0.58 -0.83 96.33 6 0.10 4.66 72.41
0.08 0.08 4.96 69.81 972 0.37 0.12 93.00 24 0.59 -0.84 95.50 4 0.09 4.82 70.47
0.12 0.07 5.27 65.31 984 0.38 0.06 93.43 14 0.60 -0.96 93.00 2 0.08 5.19 65.76
0.18 0.05 5.78 56.24 991 0.35 0.20 88.25 8 0.49 -0.74 88.00 1 0.06 5.73 56.53
0.26 0.04 6.30 42.35 998 0.42 -0.11 89.00 2 0 0.04 6.28 42.44
0.38 0.03 6.73 24.49 1,000 0 0 0.03 6.73 24.49
0.70 0.02 6.96 4.78 1,000 0 0 0.02 6.96 4.78
1.00 0.02 6.98 1.45 1,000 0 0 0.02 6.98 1.45
TABLE XV: Multi-category evaluation of samples transformed with image stylization on models trained with different distributions by using statistical metrics across object categories.
Airplane Bicycle Car
Training Distribution $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat
ImageNet1k [63] 0.02 0.01 6.94 18.00 🟊$\bigstar$ 0.03 0.02 6.84 46.00 🟊$\bigstar$ 0.04 0.03 6.42 30.00 🟊$\bigstar$
Stylised ImageNet [67] 0.55 0.27 -0.40 94.00 $\clubsuit$ 0.33 0.20 0.40 94.00 $\clubsuit$ 0.18 0.15 1.85 88.00 🟊$\bigstar$
ImageNet1k [63] + Stylised [67] 0.51 0.26 -0.33 94.00 $\clubsuit$ 0.35 0.22 0.32 96.00 $\clubsuit$ 0.15 0.12 2.54 96.00 🟊$\bigstar$
Dog Elephant Truck
Training Distribution $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Acc Pat
ImageNet1k [63] 0.02 0.01 6.95 34.00 🟊$\bigstar$ 0.02 0.01 6.98 46.00 🟊$\bigstar$ 0.02 0.01 6.91 32.00 🟊$\bigstar$
Stylised [67] 0.16 0.11 2.85 94.00 🟊$\bigstar$ 0.44 0.24 0.21 98.00 $\clubsuit$ 0.09 0.09 4.16 84.00 🟊$\bigstar$
ImageNet1k [63] + Stylised [67] 0.14 0.09 3.69 96.00 🟊$\bigstar$ 0.44 0.21 -0.26 96.00 $\clubsuit$ 0.11 0.09 3.68 82.00 🟊$\bigstar$
TABLE XVI: Performance by samples transformed with image stylization on models trained with different distributions across varied visual patterns.
Average of Classes with Pattern A (🟊$\bigstar$) (Fractured) Average of Classes with Pattern B ($\clubsuit$) (Overlapped) Average of Classes with Pattern C ($\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
Training Distribution $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc
ImageNet1k [63] 0.02 6.91 20.18 1,000 0 0 0.02 6.91 20.18
Stylised [67] 0.04 6.25 54.04 999 0.29 0.45 86.00 1 0 0.04 6.24 54.07
ImageNet1k [63] + Stylised [67] 0.04 6.41 47.97 1,000 0 0 0.04 6.41 47.97
The learned latent space of neural networks is known to be highly susceptible to slight perturbations in input samples, making it brittle. We analyze the latent space of ResNet-50 using samples modified with various input transformations to comprehend the distribution of such transformed samples. Specifically, we apply the five transformations described below to the input sample and individually evaluate their impact. Each sample is transformed once using the transformation to estimate their distribution in the latent space.
Image Crop:
The input sample is center-cropped to a size $s$ in this transformation. The reduced image is then resized to the standard $256\times 256$ dimensions and processed by the neural network. The results of this transformation are presented in Tables IX and X. Notably, as more pixels are cropped from the samples, the distribution of transformed samples becomes more fractured in the latent space. This suggests that the model relies on detecting multiple features within the image, and missing information can lead the model to perceive the sample as entirely different from the original.
Image Rotation:
[…]
Gaussian Noise
: In this transformation, Gaussian noise is added to the input sample. Mathematically, if $x$ is the input sample, $z\sim{\mathcal{N}}(\mu,\alpha^{2})$ is sampled Gaussian noise, and $\alpha$ is the strength of Gaussian Noise, transformed image $x^{\prime}$ can be written as, $x^{\prime}=x+(\alpha\times z)$, which is then processed by the neural network, and the results of this transformation are presented in Tables XIII and XIV.
Similar to the other two input transformations, this also induces fracturing of the distribution of transformed samples in the latent space. Additionally, it is noteworthy that as the strength of the Gaussian noise gradually increases, more classes become fractured, suggesting a gradual breakdown in the features identified by the model.
Image Stylization:
In this transformation, the input sample is deprived of its original texture and replaced with a random painting style [67]. The results of this transformation on models trained on the standard ImageNet-1k [63] dataset, a stylized version Stylized ImageNet [67], and a combination of both are displayed in Tables XV and XVI. Similar to other transformations, stylization in the images induces the fracturing of transformed samples. This further underscores the model’s sensitivity to variations in input samples, leading to distinct representations for stylized samples.
TABLE XVII: Multi-category evaluation of adversarial samples created with PGD Attack by using statistical metrics across object categories.
Airplane Bicycle Car
Architecture
$\epsilon$
$\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat
ResNet-50 Adversarially Trained [68] 2/255 0.33 0.22 -0.08 $\clubsuit$ 0.10 0.08 4.29 🟊$\bigstar$ 0.26 0.19 0.74 🟊$\bigstar$
4/255 0.31 0.22 0.03 $\clubsuit$ 0.09 0.08 4.22 🟊$\bigstar$ 0.17 0.13 2.00 🟊$\bigstar$
8/255 0.26 0.20 0.34 $\clubsuit$ 0.06 0.06 5.65 🟊$\bigstar$ 0.10 0.07 4.97 🟊$\bigstar$
ResNet-50 Standard Trained [58] 2/255 0.04 0.03 7.69 🟊$\bigstar$ 0.02 0.02 8.61 🟊$\bigstar$ 0.18 0.15 1.22 🟊$\bigstar$
4/255 0.06 0.06 5.83 🟊$\bigstar$ 0.02 0.02 8.60 🟊$\bigstar$ 0.29 0.22 0.10 $\clubsuit$
8/255 0.05 0.04 6.94 🟊$\bigstar$ 0.03 0.02 8.37 🟊$\bigstar$ 0.36 0.27 -0.30 $\clubsuit$
Dog Elephant Truck
Architecture
$\epsilon$
$\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat
ResNet-50 Adversarially Trained [68] 2/255 0.04 0.04 7.43 🟊$\bigstar$ 0.19 0.17 1.23 🟊$\bigstar$ 0.17 0.14 1.56 🟊$\bigstar$
4/255 0.03 0.03 7.96 🟊$\bigstar$ 0.10 0.10 3.29 🟊$\bigstar$ 0.12 0.12 2.51 🟊$\bigstar$
8/255 0.02 0.02 8.57 🟊$\bigstar$ 0.06 0.05 6.26 🟊$\bigstar$ 0.07 0.08 4.80 🟊$\bigstar$
ResNet-50 Standard Trained [58] 2/255 0.02 0.01 8.79 🟊$\bigstar$ 0.04 0.04 7.38 🟊$\bigstar$ 0.03 0.02 8.35 🟊$\bigstar$
4/255 0.02 0.02 8.60 🟊$\bigstar$ 0.05 0.05 6.37 🟊$\bigstar$ 0.04 0.04 7.54 🟊$\bigstar$
8/255 0.04 0.04 7.42 🟊$\bigstar$ 0.05 0.05 6.61 🟊$\bigstar$ 0.04 0.04 7.53 🟊$\bigstar$
TABLE XVIII: Performance on adversarial samples created with PGD Attack across varied visual patterns.
Average of Classes with Pattern A (🟊$\bigstar$) (Fractured) Average of Classes with Pattern B ($\clubsuit$) (Overlapped) Average of Classes with Pattern C ($\spadesuit$) (Clustered) Average of 1,000 ImageNet1k Classes
Architecture
$\epsilon$
$\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc
ResNet-50 Adversarially Trained [68] 2/255 0.05 6.14 50.61 983 0.34 0.05 90.43 14 0.57 -1.18 91.33 3 0.05 6.03 51.29
4/255 0.04 6.38 36.18 990 0.32 0.14 85.50 8 0.57 -1.17 88.00 2 0.04 6.31 36.68
8/255 0.03 6.64 16.45 996 0.30 0.24 72.67 3 0.50 -1.00 90.00 1 0.03 6.62 16.69
ResNet-50 Standard Trained [58] 2/255 0.05 5.82 0.00 980 0.32 0.01 0.00 15 0.50 -0.80 0.00 5 0.06 5.70 0.00
4/255 0.05 5.93 0.00 986 0.33 -0.01 0.00 11 0.50 -0.70 0.00 3 0.05 5.85 0.00
8/255 0.04 6.38 0.03 996 0.31 0.25 0.00 4 0 0.04 6.36 0.03
TABLE XIX: Multi-category evaluation using different distance metrics by using statistical metrics across object categories.
Airplane Bicycle Car
Distance Metric $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat
Euclidean ($l_{2}$ norm) 0.57 0.25 -0.84 $\spadesuit$ 0.15 0.11 2.60 🟊$\bigstar$ 0.45 0.26 -0.17 $\clubsuit$
CityBlock ($l_{1}$ norm) 0.55 0.25 -0.71 $\spadesuit$ 0.14 0.10 2.99 🟊$\bigstar$ 0.41 0.24 -0.05 $\clubsuit$
Max Norm ($l_{\infty}$ norm) 0.69 0.31 -1.02 $\spadesuit$ 0.31 0.28 0.74 $\clubsuit$ 0.42 0.26 -0.26 $\clubsuit$
Cosine 0.65 0.29 -0.96 $\spadesuit$ 0.23 0.16 1.29 🟊$\bigstar$ 0.45 0.28 -0.05 $\clubsuit$
Dog Elephant Truck
Distance Metric $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat $\mu_{k^{*}}$ $\sigma_{k^{*}}$ $\gamma_{k^{*}}$ Pat
Euclidean ($l_{2}$ norm) 0.14 0.11 2.88 🟊$\bigstar$ 0.62 0.29 -0.99 $\spadesuit$ 0.44 0.25 -0.20 $\clubsuit$
CityBlock ($l_{1}$ norm) 0.13 0.10 3.12 🟊$\bigstar$ 0.56 0.27 -0.79 $\clubsuit$ 0.39 0.23 0.05 $\clubsuit$
Max Norm ($l_{\infty}$ norm) 0.10 0.07 4.74 🟊$\bigstar$ 0.63 0.28 -1.09 $\spadesuit$ 0.53 0.30 -0.37 $\clubsuit$
Cosine 0.17 0.12 2.11 🟊$\bigstar$ 0.57 0.29 -0.46 $\spadesuit$ 0.52 0.28 -0.43 $\clubsuit$
Adversarial Perturbation:
To measure the distribution of adversarial samples, we perturb the input sample with adversarial perturbations optimized by a PGD attack [66] with varying strength $\epsilon$ of adversarial perturbation. For a given input sample $x$ and a neural network $f$ such that $f(x)$ is the label of sample $x$ predicted by the neural network, adversarial perturbation $\delta$ can be defined as, $f(x)\neq f(x+\delta)$, where, the adversarial perturbation $\delta$ can be further optimized as,
[…]
The results of adversarial samples on robust and non-robust models are presented in Tables XVII and XVIII. Like other transformations, adversarial perturbations cause also fragmentation in the distribution of samples in the latent space. Although models trained with adversarial techniques exhibit increased robustness and improved transferability as demonstrated by Kotyan et al. [71], they are not completely robust. This is indicated by a gradual rise in the fragmentation with stronger perturbations, making a evident trade-off between robustness and accuracy as highlighted by [70, 72].
In summary, these findings highlight the neural network’s sensitivity to variations in input samples, showcasing that a model perceives transformed samples as substantially different from their original counterparts. The results across various input transformations, including adversarial perturbations, indicate that current models struggle to cluster transformed samples together, interpreting them as distinct from the original sample and each other. Observations also indicate that white-box attacks such as PGD exploit this struggle by further fragmenting the distribution of samples into smaller fractures.
IV-G Effect of Different Distance Metrics on k* Distribution
To evaluate the sensitivity of the nearest neighbor method to different distance metrics and understand their impact on k* distribution and values, we compute the k* distribution using Euclidean ($l_{2}$ norm), City Block ($l_{1}$ norm), Max Norm ($l_{\infty}$ norm), and Cosine Distances. The results from Table XIX reveal the responsiveness of the k* distribution in terms of metrics such as $\mu_{k^{*}}$, $\sigma_{k^{*}}$, and $\gamma_{k^{*}}$. Despite variations, substantial agreement exists on the sample distribution classification across different distance metrics. Having said that, this trend shifts when collectively assessing performance, as shown in Table XX. Here, we observe minimal sensitivity in metric values like $\mu_{k^{*}}$ and $\gamma_{k^{*}}$ to choose distance metrics. In other words, the selection of distance metrics does impact the number of classes classified into multiple patterns.
TABLE XX: Performance using Different Distance Metrics across varied visual patterns.
Average of Classes with Pattern A (🟊$\bigstar$)  (Fractured) Average of Classes with Pattern B ($\clubsuit$)  (Overlapped) Average of Classes with Pattern C ($\spadesuit$)  (Clustered) Average of 1,000 ImageNet1k Classes
Distance Metric $\mu_{k^{*}}$ $\gamma_{k^{*}}$ N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ N $\mu_{k^{*}}$ $\gamma_{k^{*}}$
Euclidean ($l_{2}$ norm) 0.10 4.51 935 0.37 0.14 56 0.61 -1.04 9 0.12 4.21
CityBlock ($l_{1}$ norm) 0.09 4.59 949 0.37 0.15 42 0.58 -0.88 9 0.11 4.35
Max Norm ($l_{\infty}$ norm) 0.10 4.27 901 0.39 0.10 82 0.63 -1.12 17 0.14 3.83
Cosine 0.10 4.30 915 0.38 0.07 73 0.62 -1.14 12 0.13 3.93
TABLE XXI: Performance using different tasks and architectures across varied visual patterns
Average of Classes with Pattern A (🟊$\bigstar$) (Fractured) Average of Classes with Pattern B ($\clubsuit$) (Overlapped) Average of Classes with Pattern C ($\spadesuit$) (Clustered) Average on Entire Dataset
Task Architecture $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N $\mu_{k^{*}}$ $\gamma_{k^{*}}$ Acc N
Intent Classification on [64] (Text) DeBERTa V3 [73] 0.18 2.68 64.64 24 0.38 -0.05 81.12 14 0.65 -1.29 83.73 21 0.40 0.62 75.34 59
XLM-RoBERTa [73] 0.21 2.04 79.26 22 0.40 0.07 64.00 8 0.69 -1.60 90.04 29 0.47 -0.02 82.49 59
BERT [73] 0.22 2.18 64.81 15 0.40 -0.09 87.29 8 0.71 -1.75 92.38 36 0.54 -0.53 84.68 59
Multilingual-MiniLM [73] 0.20 2.17 80.22 15 0.42 -0.14 82.87 11 0.71 -1.69 91.33 33 0.53 -0.42 86.93 59
Keyword Spotting on [65] (Audio) AST [74] 0.26 1.36 70.00 1 0.43 0.00 58.83 3 0.77 -1.89 95.18 32 0.73 -1.65 91.45 36
Wav2Vec2-Conformer-L [75] 0.22 1.82 92.00 1 0.38 0.01 40.00 2 0.92 -3.81 98.04 33 0.87 -3.44 94.65 36
IV-H Going Beyond Image Classification
Exploring beyond the domain of computer vision, we choose to apply the k* distribution to models trained for tasks other than image classification. Specifically, we evaluate models trained for classifying intent from MASSIVE dataset [64] in the domain of natural language processing, and models trained for keyword spotting from the Speech Commands dataset [65] in the domain of speech processing as shown in Table XXI. We note the wider applicability of k* distribution to multiple domains and observe a similar trend of higher $\mu_{k^{*}}$ and lower $\gamma_{k^{*}}$ for better accuracy. Interestingly, we observe that the models trained for intent classification and keyword spotting were less fractured than image classification, suggesting that neural networks’ feature and label association for such tasks can be further improved.
V Conclusion
In this article, we introduce the k* distribution, a methodology grounded in the local neighborhood, to assess the distribution of samples in the learned latent space of neural networks. Our experimental findings indicate that the distribution can be primarily categorized into three distinct patterns: Pattern A (🟊$\bigstar$) representing Fractured distribution of samples in latent space: Identified by a positively skewed k* distribution. Pattern B ($\clubsuit$) representing Overlapped distribution of samples in latent space: Characterized by a nearly uniform k* distribution. Pattern C ($\spadesuit$) representing Clustered distribution of samples in latent space: Indicated by a negatively skewed k* distribution. Using k* distribution, we analyzed the learned latent space of different models, with different training data distributions, and of different layers. Further, the learned latent space was tested against various input transformations to understand how the input transformations affect the learned latent space and are associated with the training distribution. We hope this methodology and analysis will help understand the neural networks’ latent spaces and improve the distributions of samples in latent space.
Acknowledgments
This work was supported by JSPS Grant-in-Aid for Challenging Exploratory Research - Grant Number JP22K19814, JST Strategic Basic Research Promotion Program (AIP Accelerated Research) - Grant Number JP22584686, JSPS Research on Academic Transformation Areas (A) - Grant Number JP22H05194.